### v1

# Reinforced In-Context Black-Box Optimization

Abstract
Black-Box Optimization (BBO) has found successful applications in many fields of science and engineering. Recently, there has been a growing interest in meta-learning particular components of BBO algorithms to speed up optimization and get rid of tedious hand-crafted heuristics. As an extension, learning the entire algorithm from data requires the least labor from experts and can provide the most flexibility. In this paper, we propose RIBBO, a method to reinforce-learn a BBO algorithm from offline data in an end-to-end fashion. RIBBO employs expressive sequence models to learn the optimization histories produced by multiple behavior algorithms and tasks, leveraging the in-context learning ability of large models to extract task information and make decisions accordingly. Central to our method is to augment the optimization histories with regret-to-go tokens, which are designed to represent the performance of an algorithm based on cumulative regret of the histories. The integration of regret-to-go tokens enables RIBBO to automatically generate sequences of query points that satisfy the user-desired regret, which is verified by its universally good empirical performance on diverse problems, including BBOB functions, hyper-parameter optimization and robot control problems.
Machine Learning, ICML
1 Introduction
Black-Box Optimization (BBO) [4; 1] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available. To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost. Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [41; 52], material design [21; 23], robot control [10; 11], and optimal experimental design [24; 43], just to name a few.
To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [50; 20] and Evolutionary Algorithms (EA) [5; 65]. Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics. Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3; 6].
Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [18; 2]. Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization. For example, some components in BO are proposed to be learned from data, including the surrogate model [45; 58; 61; 40], acquisition function [56; 29], initialization strategy [17; 46], and search space [44; 36]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [33] or the update rules for evolution strategy [34].
There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems. However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13; 53] or online sampling from the objective function (often very expensive) [38]Chen et al. [14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge. Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion. RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing multiple behavior algorithms on multiple tasks. The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point. During testing, the sequence model itself serves as a BBO algorithm by generating the query points auto-regressively. Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm. By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret. Such modeling enables RIBBO to circumvent the impact of inferior data and further reinforce its performance on top of the behavior algorithms.
We perform experiments on BBOB synthetic functions, hyper-parameter optimization and robot control problems by using some representatives of heuristic search, EA, and BO as behavior algorithms to generate the offline datasets. The results show that RIBBO can automatically generate sequences of query points satisfying the user-desired regret across diverse problems, and achieve good performance universally. Note that the best behavior algorithm depends on the problem at hand, and RIBBO can perform even better on some problems. Compared to the most related method OptFormer [14], RIBBO also has clear advantage. In addition, we perform a series of experiments to analyze the influence of important components of RIBBO.
2 Background
2.1 Black-Box Optimization
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ be a black-box function, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional search space. The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value. Several classes of BBO algorithms have been proposed, e.g., BO [50; 20] and EA [5; 65]. The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [48], and an acquisition function [60], which are used to model $f$ and decide the next query point, respectively. EA is a class of heuristic optimization algorithms inspired by natural evolution. It maintains a population of solutions and iterates through mutation, crossover, and selection operations to find better solutions.
To evaluate the performance of BBO algorithms, regrets are often used. The instantaneous regret $r_{t}=f(\bm{x}^{*})-f(\bm{x}_{t})$ measures the gap of function values between an optimal solution $\bm{x}^{*}$ and the currently selected point $\bm{x}_{t}$. The cumulative regret ${\rm Reg}_{T}=\sum_{i=1}^{T}r_{i}$ is the sum of instantaneous regrets in the first $T$ iterations.
2.2 Meta-Learning in Black-Box Optimization
Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming. One solution is meta-learning [55; 28], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks. By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
Meta-learning particular components has been studied with different BBO algorithms. Meta-learning in BO can be divided into four main categories according to “what to transfer” [6], including the design of the surrogate model, acquisition function, initialization strategy, and search space. For surrogate model design, Wang et al. [58]; Wistuba & Grabocka [61] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al. [45]; Müller et al. [40] substituted GP with Bayesian linear regression or neural process [22; 39]. For acquisition function design, MetaBO [56] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19]. The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17; 46] or shrink the search space [44; 36]. Meta-learning in EA usually focuses on specific evolutionary operations. For example, Lange et al. [33] substituted core genetic operators, including selection and mutation rate adaptation, with dot-product attention modules; Lange et al. [34] meta-learned a self-attention-based architecture to discover effective and order-invariant update rules.
Meta-learning entire algorithms has also been explored to obtain more flexible models. Early works [13; 53] uses Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations. RNN uses its memory state to store information about history and outputs the next query point. This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO problems. OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems. It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems. However, when using OptFormer to solve problems, a user must specify an algorithm manually. How to select suitable algorithms automatically is still a challenge. Neural Acquisition Processes (NAP) [38] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly. Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency. Compared to these two state-of-the-art E2E methods, i.e., OptFormer [14] and NAP [38], our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use during the test phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
2.3 Decision Transformer
Transformer has emerged as a powerful architecture for sequence modeling tasks [62; 32; 59]. A basic building block behind transformer is the self-attention mechanism [54], which captures the correlation between tokens of any pair of timesteps. As the scale of data and model increases, transformer has demonstrated the in-context learning ability [9], which refers to the capability of the model to infer the tasks at hand based on the input contexts.
Decision Transformer (DT) [12] abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions. Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return. Inspired by DT, we will treat BBO tasks as a sequence modeling problem naturally, use a causal transformer for modeling, and train it by conditioning on future regrets. Such design is expected to enable the learned model to distinguish algorithms with different performance and achieve good performance with a user-specified low regret.
(a) Data Generation
(b) Training and Inference
Figure 1: Illustration of RIBBO. Left: Data Generation. $K$ existing BBO algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ and $N$ BBO tasks $\{f_{i}\}_{i=1}^{N}$ are used to serve as the behavior algorithms and the training tasks, respectively. The offline datasets $\{\mathcal{D}_{i,j}\}$ consist of the optimization histories $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ collected by executing each behavior algorithm $\mathcal{A}_{j}$ on each task $f_{i}$ for $T$ evaluation steps, which are then augmented with the regret-to-go tokens $R_{t}$ (calculated as the cumulative regret $\sum^{T}_{t^{\prime}=t+1}(y^{*}-y_{t^{\prime}})$ over the future optimization history) to generate the final dataset $\{\widehat{\mathcal{D}}_{i,j}\}$ for training. Right: Training and Inference. Our model takes in triplets of $(\bm{x}_{t},y_{t},R_{t})$, embeds them into one token, and outputs the distribution over the next query point $\bm{x}_{t+1}$. During training, the ground-truth next query is used to minimize the loss in Eq. (4). During inference, the next query point $\bm{x}_{t+1}$ is generated auto-regressively based on the current history $\hat{\bm{h}}_{t}$.
3 Method
This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion. The overall workflow is presented in Figure 1.
3.1 Problem Formulation
We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19; 28; 64]. Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
Let $[N]$ denote the integer set $\{1,2,\ldots,N\}$. During training, we usually access $N$ source tasks and each task corresponds to an objective function $f_{i}\sim P(\mathcal{F})$, where $i\in[N]$. Hereafter, we use $f_{i}$ to denote the task $i$ if the context is clear. We assume that the information is available via offline datasets $\mathcal{D}_{i,j}$, which are produced by executing a behavior algorithm $\mathcal{A}_{j}$ on task $f_{i}$, where $j\in[K]$ and $i\in[N]$. Each dataset $\mathcal{D}_{i,j}=\{\bm{h}^{i,j,m}_{T}\}_{m=1}^{M}$ consists of $M$ optimization histories $\bm{h}^{i,j,m}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$, where $\bm{x}_{t}$ is the query point selected by $\mathcal{A}_{j}$ at iteration $t$, and $y_{t}=f_{i}(\bm{x}_{t})$ is its objective value. If the context is clear, we will omit $i,j,m$ and simply use $\bm{h}_{T}$ to denote an optimization history with length $T$. The initial history $\bm{h}_{0}$ is defined as $\emptyset$. We impose no additional assumptions about the behavior algorithms, allowing for a range of BBO algorithms, even random search.
With the datasets, we seek to learn a model $\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})$, which is parameterized by $\bm{\theta}$ and generates the next query point $\bm{x}_{t}$ by conditioning on the previous history $\bm{h}_{t-1}$. As introduced in Section 2.1, with a given budget $T$ and the history $\bm{h}_{T}$ produced by an algorithm $\mathcal{A}$, we use the cumulative regret
${\rm Reg}_{T}=\sum_{t=1}\nolimits^{T}(y^{*}-y_{t})$ (1)
as the evaluation metric, where $y^{*}$ is the optimum value and $\{y_{t}\}^{T}_{t=1}$ are the function values of the points in $\bm{h}_{T}$.
3.2 Method Outline
Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$. Similar to the previous work [13], we take this framework as a starting point and define the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$. The optimization histories in offline datasets provide a natural supervision for learning.
Suppose we have a set of histories $\{\bm{h}_{T}\}$, generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$. By employing a causal transformer model $\mathcal{M}_{\bm{\theta}}$, we expect $\mathcal{M}_{\bm{\theta}}$ to imitate $\mathcal{A}$ and produce similar optimization history on $f$. In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$. To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
$\mathcal{L}_{\rm BC}(\bm{\theta})=-\mathbb{E}_{\bm{h}_{T}\sim\mathcal{D}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})\right].$ (2)
To effectively minimize the loss, $\mathcal{M}_{\bm{\theta}}$ needs to recognize both the task and the behavior algorithm in-context, and then imitate the optimization behavior of the corresponding behavior algorithm.
Nevertheless, naively imitating the offline datasets hinders the model since some inferior behavior algorithms may severely degenerate the model’s performance. Inspired by DT [12], we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
$\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T},\;R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}}),$ (3)
where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding, denoted as [PAD] in Figure 1(b), and $R_{T}=0$. The augmented histories compose the augmented dataset $\widehat{\mathcal{D}}_{i,j}$, and the training objective of $\mathcal{M}_{\bm{\theta}}$ becomes
$\mathcal{L}_{\rm{RIBBO}}(\bm{\theta})=-\mathbb{E}_{\hat{\bm{h}}_{T}\sim\widehat{\mathcal{D}}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\hat{\bm{h}}_{t-1})\right].$ (4)
The integration of RTG tokens offers several advantages. [label=0)] 1. RTG tokens in the context bring identifiability of behavior algorithms, and the model $\mathcal{M}_{\bm{\theta}}$ can effectively utilize them to make appropriate decisions. 2. RTG tokens have a direct correlation with the metric of interest, i.e., cumulative regret $\mathrm{Reg}_{T}$ in Eq. (1). Conditioning on a lower RTG token provides a guidance to our model and reinforces $\mathcal{M}_{\bm{\theta}}$ to exhibit superior performance. These advantages will be clearly shown by experiments in Section 4.4.
Note that the resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context. The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31]. It has been explored to infer general functional relationships as supervised learning or RL algorithms. For example, Guo et al. [25]; Hollmann et al. [27]; Li et al. [37] feed the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context and expect the model to make an accurate prediction on the query input $\bm{x}_{t}$askin et al. [35] learn RL algorithms using causal transformers. Here, we use it for BBO.
3.3 Practical Implementation
Next, we detail the model architecture, training and inference of RIBBO.
Model Architecture. For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture [47], which comprises a stack of causal attention blocks, each containing an attention component and a feed-forward network. We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network. The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point.
Note that previous works that adopt the sequence model as the surrogate model [42; 39; 40] typically remove the positional encoding because the surrogate model should be invariant to the history order. On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms. Additionally, the positional encoding can keep the monotonically decreasing order of RTG tokens. More architecture details can be found in Appendix A.
Model Training. RTG tokens are computed as outlined in Eq. (3) for the offline datasets before training. As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum value. Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms. We sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets and train $\mathcal{M}_{\bm{\theta}}$ to minimize the RIBBO loss in Eq. (4).
Algorithm 1 Model Inference with HRR
Input: trained model $\mathcal{M}_{\bm{\theta}}$, budget $T$, optimum value $y^{*}$
Process:
1:  Initialize context $\hat{\bm{h}}_{0}=\{(\bm{x}_{0},y_{0},R_{0})\}$, where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding and $R_{0}=0$;
2:  for $t=1,2,\ldots,T$ do
3:     Generate the next query point $\bm{x}_{t}\sim\mathcal{M}_{\bm{\theta}}(\bm{\cdot}|\hat{\bm{h}}_{t-1})$;
4:     Evaluate $\bm{x}_{t}$ to obtain $y_{t}=f(\bm{x}_{t})$;
5:     Calculate the instantaneous regret $r=y^{*}-y_{t}$;
6:     Relabel $R_{i}\leftarrow R_{i}+r$, for each $(\bm{x}_{i},y_{i},R_{i})$ in $\hat{\bm{h}}_{t-1}$;
7:     $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},0)\}$;
8:  end for
Model Inference. The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during the inference process, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$. A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$. Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and then decrease it as
$R_{t}=R_{t-1}-(y^{*}-y_{t}).$ (5)
However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
Given the fact that RTGs are lower bounded by 00 and a 00 value implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 0, and use a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sampling. The inference procedure with HRR is detailed in Algorithm 1. In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$. At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is first fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, and $\bm{x}_{t}$ is evaluated to obtain $y_{t}=f(\bm{x}_{t})$ in line 4. Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in the history $\hat{\bm{h}}_{t-1}$: calculating the instantaneous regret $r=y^{*}-y_{t}$ in line 5 and adding $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ in line 6, i.e.,
$\forall 0\leq i<t,R_{i}\leftarrow R_{i}+(y^{*}-y_{t}).$ (6)
Note that this relabelling process guarantees that $\forall 0\leq i<t$, the RTG token $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$, which can also be written as $\sum_{t^{\prime}=i+1}^{T}(y^{*}-y_{t^{\prime}})$, consistent with the definition in Eq. (3), because the immediate RTG $R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}})$ is set to 0. In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$. The above process is repeated until reaching the budget $T$. Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using zero as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over the future optimization history.
3.4 Data Generation
Finally, we give some guidelines about data generation for using the proposed RIBBO method.
Data Collection. Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection. For example, we can select some representatives of different types of BBO algorithms (e.g., BO and EA). Datasets $\mathcal{D}_{i,j}$ are then obtained by using each behavior algorithm to optimize each function with different random seeds. Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented using RTG tokens $R_{t}$ computed as in Eq. (3). The resulting augmented histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for training the model $\mathcal{M}_{\bm{\theta}}$.
Data Normalization. To provide a unified interface and balance the statistic scales across tasks, it is important to apply normalization to the inputs to our model. We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively. For the function value $y$, we apply random scaling akin to previous works [61; 14]. That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$. The RTG tokens are calculated accordingly with the normalized values. The random normalization can make a model exhibit invariance across various scales of $y$.
4 Experiments
In this section, we examine the performance of RIBBO on a wide range of tasks, including synthetic functions, Hyper-Parameter Optimization (HPO) and robot control problems. The model architecture and hyper-parameters are maintained consistently across these problems. We train our model using five distinct random seeds, ranging from 00 through $4$, and each trained model is also run five times independently during the execution phase. We will report the average performance and standard deviation. Details of the model hyper-parameters can be found in Appendix A. Our code is available at https://github.com/songlei00/RIBBO.
4.1 Experimental Setup
Benchmarks. We use BBO Benchmarks BBOB [15], HPO-B [2], and rover trajectory planning tasks [57]. The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions. For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties. According to the properties of the functions, they can be further divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems. It fits an XGBoost model as the objective function for each HPO problem, and we conduct experiments on two widely used machine models, SVM and XGBoost. For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation. Similar to [15; 56], we implement random translations and scalings to the search space to construct a distribution of functions. For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors. Detailed explanations of the benchmarks can be found in Appendix B.1.
Data. Similar to OptFormer [14], we employ seven behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [49], Eagle Strategy [63], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively. Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds. More details about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
4.2 Baselines
As our proposed method RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with an offline dataset, including Behavior Cloning (BC) [7] and OptFormer [14]. Their hyper-parameters are set as same as that of our model for fairness. Note that the seven behavior algorithms used to generate datasets are also important baselines, which are included for comparison as well.
Figure 2: Performance comparison among RIBBO, BC, BC Filter, OptFormer, and behavior algorithms on synthetic functions, HPO, and robot control problems. The $y$-axis is the normalized average objective value, and the length of vertical bars represents the standard deviation.
BC uses the same transformer architecture as RIBBO. The only difference is that we do not feed RTG tokens into the model of BC and train to minimize the BC loss in Eq. (2). When the solutions are generated auto-regressively, BC tends to imitate the average behavior of various behavior algorithms. Consequently, the inclusion of underperforming algorithms, e.g., Random Search and Shuffled Grid Search, may significantly degrade the performance. To mitigate this issue, we have also trained the model by excluding these underperforming algorithms, denoted as BC Filter.
OptFormer employs transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution. Its original implementation is built upon a text-based transformer with a large training scale. In this paper, we re-implemented a simplified version of OptFormer where we only maintain the algorithm type within the metadata, i.e., employing different initial states $\bm{x}_{0}$ and $y_{0}$ for different algorithms. More details about the training procedure and imitation capacity of the reimplementation can be found in Appendix C.
4.3 Main Results
The results are shown in Figure 2. For the sake of clarity in visualization, we have omitted the inclusion of Random Search and Shuffled Grid Search due to their poor performance from start to finish. We can observe that RIBBO achieves superior or at least equivalent efficacy in comparison to the best behavior algorithm on each problem except SVM and rover. This demonstrates the versatility of RIBBO, while the most effective behavior algorithm depends upon the specific problem at hand, e.g., the best behavior algorithms on Lunacek, Rastrigin and XGBoost are GP-EI, Eagle Strategy and CMA-ES, respectively. Note that the good performance of RIBBO does not owe to the memorization of optimal solutions, as the search space is transformed randomly, resulting in variations in optimal solutions across different functions from the same distribution. It is because RIBBO is capable of using RTG tokens to identify algorithms and reinforce the performance on top of the behavior algorithms, which will be clearly shown later. We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing zero as the immediate RTG to generate the optimal potential solutions.
RIBBO does not perform well on SVM problem, which may be due to the problem’s low-dimensional nature (only three parameters) and its relative simplicity for optimization. Behavior algorithms can easily achieve good performance, while the complexity of RIBBO’s training and inference processes could instead result in the performance degradation. Regarding to rover problem where GP-EI performs the best, we collect less data from GP-EI than other behavior algorithms due to the high time cost. This may limit RIBBO’s capacity to leverage the high-quality data from GP-EI, given its small proportion relative to the data collected from other behavior algorithms. Despite this, RIBBO is still the runner-up, significantly surpassing the other behavior algorithms.
(a) Visualization
(b) Cross-dist generalization
(c) Initial RTGs
(d) RTG update strategy
Figure 3: (a) Visualization of the contour lines of 2D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO. (b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties. Note that the legend shares with that of Figure 2. (c) Initial RTG $R_{0}$’s influence on performance. (d) RTG update strategy comparison between HRR and the naive strategy with different $R_{0}$.
Compared with BC and BC Filter, RIBBO performs consistently better except on the SVM problem. BC tends to imitate the average behavior of various algorithms, and its poor performance is due to the aggregation of data from behavior algorithms with inferior performance. BC Filter is generally better than BC, because the data from the two underperforming behavior algorithms, i.e., Random Search and Shuffled Grid Search, are manually excluded for the training of BC Filter. As introduced before, OptFormer requires to manually specify which algorithm to execute. We have specified the behavior algorithm Eagle Strategy in Figure 2, which obtains good overall performance on these problems. It can be observed that OptFormer displays a close performance to Eagle Strategy, while RIBBO performs better. More results about the imitation capacity of OptFormer can be found in Appendix C.
Why Does RIBBO Behave Well? To better understand RIBBO, we train the model using only two behavior algorithms, Eagle Strategy and Random Search, which represent a good algorithm and an underperforming one, respectively. Figure 3(a) visualizes the contour lines of the 2D Branin function and the sampling points of RIBBO, Eagle Strategy, and Random Search, represented by red, orange, and gray points, respectively. The arrows are used to represent the optimization trajectory of RIBBO. Note that the two parameters of Branin have been scaled to $[-1,1]$ for better visualization. It can be observed that RIBBO makes a prediction preferring Eagle Strategy over Random Search, indicating its capability to automatically identify the quality of training data. Additionally, RIBBO achieves the exploration and exploitation trade-off capability using its knowledge about the task obtained during training, thus generating superior solutions over the ones in the training dataset.
Cross-Distribution Generalization. We further conduct experiments to examine the generalization of RIBBO to unseen function distributions. For this purpose, we train the model on $4$ of the $5$ chosen synthetic functions and test it with the remaining one. Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and we sample a set of functions from each distribution as introduced before. Greiwank Rosenbrock is used for testing. The results in Figure 3(b) suggest that RIBBO has a strong generalizing ability to unseen function distributions. More results on cross-distribution generalization are deferred to Appendix D.
The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in [13]. In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
4.4 Ablation Studies
RIBBO augments the histories with RTG tokens, facilitating distinguishing algorithms and automatically generating algorithms with user-specified performance. Next, we will verify the effectiveness of RTG conditioning and HRR strategy.
Influence of Initial RTG Token $R_{0}$. By incorporating RTG tokens into modeling, RIBBO should be able to attend to the RTGs and generate optimization trajectories that achieve the specified initial RTG token $R_{0}$. To validate this, we examine the performance of RIBBO when specifying different values of $R_{0}$, and the results are presented in Figure 3(c). Here, the RTG values, instead of the normalized objective values, of the trajectory are used as the $y$-axis. We can observe that the cumulated regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
Effectiveness of HRR. A key point of the inference procedure is how to update the value of RTG tokens at each iteration. To assess the effectiveness of the proposed strategy HRR in Eq. (6), we compare it with the naive strategy in Eq. (5), which sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration. The results are shown in Figure 3(d). The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$. When $R_{0}=0$ i.e. the lower bound of regret, the model performs well in the initial stage. However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values. Using $R_{0}=5$ compromises the initial performance, as the model may not choose the most aggressive solutions with a high $R_{0}$. However, higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in the late stage. The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG as 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq. (6) keeps the values meaningful and feasible.
Other Ablation Studies. We also study the effect of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the length $\tau$ of sampled subsequences. Due to space limitation, we provide the results in Appendix E.
5 Conclusion
This paper proposes RIBBO, which employs a transformer architecture to learn a reinforced BBO algorithm from offline datasets in an E2E fashion. By incorporating RTG tokens into the optimization histories, RIBBO can automatically generate optimization trajectories satisfying the user-desired regret. Comprehensive experiments on BBOB, HPO and robot control problems show the versatility of RIBBO. This work is a preliminary attempt towards universal BBO, and we hope it can encourage more explorations in this direction. For example, we only consider the model training over continuous search space with the same dimensionality, and it would be interesting to explore heteroscedastic search space with different types of variables.
Potential Broader Impact
This paper presents work whose goal is to advance the field of black-box optimization. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Model Details
We employ the commonly used GPT architecture [47] and the hyper-parameters are maintained consistently across the problems. Details can be found in Table 1. For BC, BC Filter and OptFormer, the hyper-parameters are set as same as that of our model.
Table 1: List of hyper-parameter settings in RIBBO.
RIBBO
Embedding dimension 256
Number of self-attention layers 12
Number of self-attention heads 8
Point-wise feed-forward dimension 1024
Dropout rate 0.1
Batch size 64
Learning rate 0.0002
Learning rate decay 0.01
Optimizer Adam
Optimizer scheduler Linear warm up and cosine annealing
Number of training steps 500,000
Length of subsequence $\tau$ 50
Appendix B Details of Experimental Setup
B.1 Benchmarks
  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in continue domain. This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties. According to the properties of functions, these synthetic functions can be divided into $5$ categories, i.e., (1) separable functions, (2) moderately conditioned functions, (3) ill-conditioned and unimodal functions, (4) multi-modal functions with adequate global structure, and (5) multi-modal functions with weak global structures. We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5). We use the BBOB benchmark implementation in Open Source Vizier. The dimension is set to $10$ for all functions.
  • SVM and XGBoost. HPO-B [2] is the most commonly used HPO benchmark and grouped by search space id. Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost. Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset. For continue domain, it fits an XGBoost model as the objective function for each HPO problem. These datasets for each search space id are divided into training and test datasets. We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively. SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most parameters in HPO-B.
  • Rover Trajectory Planning [57; 16] is a trajectory optimization task to emulate a rover navigation task. The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, totally $60$ parameters to optimize. Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{x}_{s}\|_{1}+\|\bm{x}_{58,59}-\bm{x}_{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of a given trajectory, $\bm{x}_{s}$ and $\bm{x}_{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem. This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions. To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ as well as scalings in $[0.9,1.1]$, similar to [56; 15]. The training and test tasks are randomly sampled from the distribution.
B.2 Bahavior Algorithms
The datasets are generated with some representatives of heuristic search, EA, and BO as behavior algorithms. We use the implementation in Open Source Vizier [51] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES. For Hill Climbing and Regularized Evolution, we provide a simple reimplementation. We use the implementation in BoTorch [8] for GP-EI. The details of behavior algorithms are summarized as follows.
  • Random Search uniformly at random selects a point in the domain at each iteration.
  • Shuffled Grid Search discretizes the ranges of real parameters into $100$ equidistant points and then selects without replacement a random point from the grid at each iteration.
  • Hill Climbing. At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to $x_{next}$ and evaluated. If $f(x_{next})\textgreater f(x_{best})$, we update $x_{best}$ to $x_{next}$.
  • Regularized Evolution [49] is an evolutionary algorithm with tournament selection and age-based replacement. We use a population size of $25$ and tournament size of $5$. At each iteration, we randomly select a tournament subset from the current population, and mutate the solution with maximum value. The mutation operation is to uniformly at random selects one of the parameters and mutates to a random value in the domain.
  • Eagle Strategy [63] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies. Each firefly emits a light whose intensity corresponds to the objective value. At each iteration, for a firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker one in its vicinity, and the position is updated using the calculated weight.
  • CMA-ES [26] is a popular evolutionary algorithm. At each iteration, the candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
  • GP-EI employs GP as the surrogate model and EI [30] as the acquisition function for BO.
B.3 Data
For training, a set of tasks are sampled from the task distribution and above behavior algorithms are used to collect data. Specifically, for BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times except GP-EI, which is limited to $100$ function samples due to high time cost. For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost. All behavior algorithms employ $500$ random seeds in the execution except GP-EI, which is executed using $500$ random seeds. For rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times. For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
Appendix C Re-Implementation of OptFormer
OptFormer [14] is a general optimization framework based on transformer [54] and provides an interface to learn policy or function prediction. When provided with textual representations of historical data, OptFormer is capable of determining the next query point $\bm{x}_{t}$ as policy. Additionally, if the histories incorporate a possible query point $\bm{x}_{t}$, the framework is designed to predict the corresponding $y_{t}$, thereby serving as function prediction. We concentrate on the aspect of policy learning, attributable to its greater relevance to our work.
The original implementation is built upon text-based transformer and uses private datasets for training. In this paper, we have re-implemented a simplified version of OptFormer, where we omit the textual tokenization process and only maintain the algorithm type within the metadata. Numerical inputs are fed into the model and we employ different initial states $\bm{x}_{0}$ and $y_{0}$ to represent different algorithms. The hyper-parameters are set as same as our method.
To examine the algorithm imitation capability, we compare our re-implemented with the corresponding behavior algorithms. The results are shown in Figure 4. For the sake of clarity in the visual representation, we only plot a subset of behavior algorithms, including Shuffled Grid Search, Hill Climbing, Regularized Evolution and Eagle Strategy. These behavior algorithms are plotted by solid lines, while their OptFormer counterparts are shown by dashed lines with the same color. Note that the figure is shown using the immediate function values as the $y$-axis to facilitate a comprehensive observation of the optimization.
Figure 4: Comparison of the behavior algorithms with the OptFormer re-implementation.
Appendix D Cross-Distribution Generalization
We conduct experiments to examine the generalization of our method to unseen function distributions. We train the model on $4$ of $5$ chosen synthetic functions and test it with the remaining one. The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
Figure 5: Cross-Distribution Generalization by training on $4$ synthetic functions and testing on the remaining one.
Appendix E Ablation Studies
We provide further ablation studies to examine the influence of the important components and hyper-parameters of RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
Token Aggregation aims to aggregate the information of $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them. RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token. It is compared with two different token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave). The addition method integrates the values of each token to form a single token, while interleaving method addresses each token sequentially. The results are shown in Figure 6(a). Concatenation method surpasses both the addition and interleaving methods, with the addition method exhibiting the least efficacy. The concatenation method employs a relatively straightforward technique to aggregate each token by direct concatenation, while the interleaving method adopts a more complex process to learn the interrelations of tokens. The inferior performance of the addition method is because that the summation will omit some details of the original tokens, which are necessary for the generation of the next query $\bm{x}_{t}$.
Normalization Method is to balance the scales of function value $y$ across tasks. We compare the employed random normalization with the dataset normalization and no normalization. Dataset normalization scales the value $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values for $f_{i}$, while no normalization entails utilizing the original function values. The results are shown in Figure 6(b). Random normalization and dataset normalization have similar performance, while no normalization leads to the poor performance. The normalization method has a direct effect on the calculation of RTG tokens and thereby influences the generating with a desired regret. No normalization results in significant variances in the scale of $y$, posing a challenge for the generation process. Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating the training and inference. Nevertheless, random normalization can make a model exhibit invariance across various scales of $y$ as mentioned in [61; 14]. Consequently, we recommend to use random normalization in practice.
Model Size has an effect on the in-context learning ability. We compare the performance of the employed model size with a smaller one (8 layers, 4 attention heads and 128-dimensional embedding space) and a larger one (16 layers, 12 attention heads and 384-dimensional embedding space). The results are shown in Figure 6(c), indicating that the model’s size has small impact on its overall performance. A smaller model, with limited capacity, exhibits a diminished capability to the in-context learning ability, while a larger model requires more training data, potentially leading to a slight decrement in performance.
Subsequence Length $\tau$ controls the length of context during both training and inference phases. Employing subsequence sampling can be viewed as a form of data augmentation that facilitates the training process. As shown in Figure 6(d), sampling subsequences instead of using the entire history as context, specifically when $\tau=T=150$, leads to an improvement of the performance. Causal transformer training and inference are quadratic in the context length, hence, a smaller $\tau$ can result in reduced computational complexity. However, an excessively shortened context may contain insufficient historical information, which could, in turn, adversely affect the performance.
(a) Token aggregation
(b) Normalization method
(c) Model size
(d) Subsequence length $\tau$
Figure 6: Ablation studies of token aggregation, normalization method, model size, and the subsequence length $\tau$.

### v2

@@
 RIBBO employs expressive sequence models to learn the optimization histories produced by multiple behavior algorithms and tasks, leveraging the in-context learning ability of large models to extract task information and make decisions accordingly.
-Central to our method is to augment the optimization histories with regret-to-go tokens, which are designed to represent the performance of an algorithm based on cumulative regret of the histories.
-The integration of regret-to-go tokens enables RIBBO to automatically generate sequences of query points that satisfy the user-desired regret, which is verified by its universally good empirical performance on diverse problems, including BBOB functions, hyper-parameter optimization and robot control problems.
-Machine Learning, ICML
+Central to our method is to augment the optimization histories with regret-to-go tokens, which are designed to represent the performance of an algorithm based on cumulative regret over the future part of the histories.
+The integration of regret-to-go tokens enables RIBBO to automatically generate sequences of query points that satisfy the user-desired regret, which is verified by its universally good empirical performance on diverse problems, including BBO benchmark functions, hyper-parameter optimization and robot control problems.
 1 Introduction
-Black-Box Optimization (BBO) [4; 1] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available.
+Black-Box Optimization (BBO) [1, 4] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available.
 To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost.
-Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [41; 52], material design [21; 23], robot control [10; 11], and optimal experimental design [24; 43], just to name a few.
-To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [50; 20] and Evolutionary Algorithms (EA) [5; 65].
+Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [42, 55], material design [21, 23], robot control [10, 11], and optimal experimental design [24, 44], just to name a few.
+To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [20, 51] and Evolutionary Algorithms (EA) [5, 68].
 Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics.
-Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3; 6].
-Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [18; 2].
+Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3, 6].
+Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [2, 18].
 Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization.
-For example, some components in BO are proposed to be learned from data, including the surrogate model [45; 58; 61; 40], acquisition function [56; 29], initialization strategy [17; 46], and search space [44; 36]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [33] or the update rules for evolution strategy [34].
+For example, some components in BO are proposed to be learned from data, including the surrogate model [41, 46, 61, 64], acquisition function [29, 59], initialization strategy [17, 47], and search space [37, 45]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [34] or the update rules for evolution strategy [35].
 There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems.
-However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13; 53] or online sampling from the objective function (often very expensive) [38]Chen et al.
+However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13, 56] or online sampling from the objective function (often very expensive) [39]Chen et al.
 [14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge.
@@
 In this paper, we introduce Reinforced In-context BBO (RIBBO), which learns a reinforced BBO algorithm from offline datasets in an E2E fashion.
-RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing multiple behavior algorithms on multiple tasks.
+RIBBO employs an expressive sequence model, i.e., causal transformer, to fit the optimization histories in the offline datasets generated by executing diverse behavior algorithms on multiple tasks.
 The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point.
@@
 Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm.
+A novel Hindsight Regret Relabelling (HRR) strategy is proposed to update the RTG tokens during testing.
 By integrating the RTG tokens into the modeling, RIBBO can automatically identify different algorithms, and generate sequences of query points that satisfy the specified regret.
@@
 The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value.
-Several classes of BBO algorithms have been proposed, e.g., BO [50; 20] and EA [5; 65].
+Several classes of BBO algorithms have been proposed, e.g., BO [20, 51] and EA [5, 68].
 The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [49], and an acquisition function [63], which are used to model $f$ and decide the next query point, respectively.
@@
 Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming.
-One solution is meta-learning [55; 28], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks.
+One solution is meta-learning [28, 58], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks.
 By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
@@
 For surrogate model design, Wang et al.
-[58]; Wistuba & Grabocka [61] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al.
-[45]; Müller et al.
-[40] substituted GP with Bayesian linear regression or neural process [22; 39].
+[61] and Wistuba and Grabocka [64] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al.
+[46] and Müller et al.
+[41] substituted GP with Bayesian linear regression or neural process [22, 40].
 For acquisition function design, MetaBO [59] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19].
-The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17; 46] or shrink the search space [44; 36].
+The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17, 47] or shrink the search space [37, 45].
 Meta-learning in EA usually focuses on specific evolutionary operations.
-For example, Lange et al.
-[33] substituted core genetic operators, including selection and mutation rate adaptation, with dot-product attention modules; Lange et al.
-[34] meta-learned a self-attention-based architecture to discover effective and order-invariant update rules.
+For example, Lang et al.
+substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules [34], and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules [35].
 Meta-learning entire algorithms has also been explored to obtain more flexible models.
-Early works [13; 53] uses Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations.
+Early works [13, 56] use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations.
 RNN uses its memory state to store information about history and outputs the next query point.
-This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO problems.
+This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO.
 OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems.
-It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems.
-However, when using OptFormer to solve problems, a user must specify an algorithm manually.
-How to select suitable algorithms automatically is still a challenge.
+It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems, which, however, requires a user to specify an algorithm manually when applying it to solve problems.
+Thus, how to select suitable algorithms automatically is still a challenge.
 Neural Acquisition Processes (NAP) [39] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly.
 Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency.
-Compared to these two state-of-the-art E2E methods, i.e., OptFormer [14] and NAP [38], our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use during the test phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
+Black-box Optimization NETworks (BONET) [33] employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing.
+Compared to the above state-of-the-art E2E methods, we consider the meta-BBO setting, where the training datasets consist of diverse algorithms across different functions.
+Our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
 2.3 Decision Transformer
-Transformer has emerged as a powerful architecture for sequence modeling tasks [62; 32; 59].
+Transformer has emerged as a powerful architecture for sequence modeling tasks [32, 62, 65].
 A basic building block behind transformer is the self-attention mechanism [57], which captures the correlation between tokens of any pair of timesteps.
@@
 Our model takes in triplets of $(\bm{x}_{t},y_{t},R_{t})$, embeds them into one token, and outputs the distribution over the next query point $\bm{x}_{t+1}$.
-During training, the ground-truth next query is used to minimize the loss in Eq.
+During training, the ground-truth next query point is used to minimize the loss in Eq.
 (4).
@@
 3 Method
-This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion. The overall workflow is presented in Figure 1.
-3.1 Problem Formulation
-We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19; 28; 64].
+This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion, as illustrated in Figure 1.
+We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19, 28, 67].
 Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
@@
 Each dataset $\mathcal{D}_{i,j}=\{\bm{h}^{i,j,m}_{T}\}_{m=1}^{M}$ consists of $M$ optimization histories $\bm{h}^{i,j,m}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$, where $\bm{x}_{t}$ is the query point selected by $\mathcal{A}_{j}$ at iteration $t$, and $y_{t}=f_{i}(\bm{x}_{t})$ is its objective value.
-If the context is clear, we will omit $i,j,m$ and simply use $\bm{h}_{T}$ to denote an optimization history with length $T$.
+If the context is clear, we will omit $i,j,m$ and simply use $\bm{h}_{T}$ to denote a history with length $T$.
 The initial history $\bm{h}_{0}$ is defined as $\emptyset$.
@@
 ${\rm Reg}_{T}=\sum_{t=1}\nolimits^{T}(y^{*}-y_{t})$ (1)
-as the evaluation metric, where $y^{*}$ is the optimum value and $\{y_{t}\}^{T}_{t=1}$ are the function values of the points in $\bm{h}_{T}$.
+as the evaluation metric, where $y^{*}$ is the optimum value and $\{y_{t}\}^{T}_{t=1}$ are the function values in $\bm{h}_{T}$.
 3.1 Method Outline
 Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$.
-Similar to the previous work [13], we take this framework as a starting point and define the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$.
+Similar to the previous work [13], we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$.
 The optimization histories in offline datasets provide a natural supervision for learning.
-Suppose we have a set of histories $\{\bm{h}_{T}\}$, generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$.
-By employing a causal transformer model $\mathcal{M}_{\bm{\theta}}$, we expect $\mathcal{M}_{\bm{\theta}}$ to imitate $\mathcal{A}$ and produce similar optimization history on $f$.
+Given a set of histories $\{\bm{h}_{T}\}$ generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$, we employ a causal transformer model $\mathcal{M}_{\bm{\theta}}$, and expect it to imitate $\mathcal{A}$ and produce similar optimization history on $f$.
 In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$.
 To fit $\mathcal{M}_{\bm{\theta}}$, we use the negative log-likelihood loss
-$\mathcal{L}_{\rm BC}(\bm{\theta})=-\mathbb{E}_{\bm{h}_{T}\sim\mathcal{D}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})\right].$ (2)
+$\mathcal{L}_{\rm BC}(\bm{\theta})=-\mathbb{E}_{\bm{h}_{T}\sim\mathcal{D}_{i,j}}\left[\sum\nolimits_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\bm{h}_{t-1})\right].$ (2)
 To effectively minimize the loss, $\mathcal{M}_{\bm{\theta}}$ needs to recognize both the task and the behavior algorithm in-context, and then imitate the optimization behavior of the corresponding behavior algorithm.
@@
 Inspired by DT [12], we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
-$\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T},\;R_{t}=\sum_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}}),$ (3)
+$\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T},\;R_{t}=\sum\nolimits_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}}),$ (3)
 where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding, denoted as [PAD] in Figure 1(b), and $R_{T}=0$.
 The augmented histories compose the augmented dataset $\widehat{\mathcal{D}}_{i,j}$, and the training objective of $\mathcal{M}_{\bm{\theta}}$ becomes
-$\mathcal{L}_{\rm{RIBBO}}(\bm{\theta})=-\mathbb{E}_{\hat{\bm{h}}_{T}\sim\widehat{\mathcal{D}}_{i,j}}\left[\sum_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\hat{\bm{h}}_{t-1})\right].$ (4)
-The integration of RTG tokens offers several advantages.
-[label=0)] 1.
-RTG tokens in the context bring identifiability of behavior algorithms, and the model $\mathcal{M}_{\bm{\theta}}$ can effectively utilize them to make appropriate decisions.
-2.
-RTG tokens have a direct correlation with the metric of interest, i.e., cumulative regret $\mathrm{Reg}_{T}$ in Eq.
+$\mathcal{L}_{\rm{RIBBO}}(\bm{\theta})=-\mathbb{E}_{\hat{\bm{h}}_{T}\sim\widehat{\mathcal{D}}_{i,j}}\left[\sum\nolimits_{t=1}^{T}\log\mathcal{M}_{\bm{\theta}}(\bm{x}_{t}|\hat{\bm{h}}_{t-1})\right].$ (4)
+The integration of RTG tokens in the context brings identifiability of behavior algorithms, and the model $\mathcal{M}_{\bm{\theta}}$ can effectively utilize them to make appropriate decisions.
+Furthermore, RTG tokens have a direct correlation with the metric of interest, i.e., cumulative regret $\mathrm{Reg}_{T}$ in Eq.
 (1).
@@
 These advantages will be clearly shown by experiments in Section 4.4.
-Note that the resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context.
+The resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context.
 The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31].
 It has been explored to infer general functional relationships as supervised learning or RL algorithms.
-For example, Guo et al.
-[25]; Hollmann et al.
-[27]; Li et al.
-[37] feed the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context and expect the model to make an accurate prediction on the query input $\bm{x}_{t}$askin et al.
-[35] learn RL algorithms using causal transformers.
+For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context [25, 27, 38]Laskin et al.
+[36] learned RL algorithms using causal transformers.
 Here, we use it for BBO.
 3.2 Practical Implementation
-Next, we detail the model architecture, training and inference of RIBBO.
+Next, we detail the model architecture, training, and inference of RIBBO.
 Model Architecture.
@@
 The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point.
-Note that previous works that adopt the sequence model as the surrogate model [42; 39; 40] typically remove the positional encoding because the surrogate model should be invariant to the history order.
+Note that previous works that adopt the sequence model as a surrogate model [40, 41, 43] typically remove the positional encoding because the surrogate model should be invariant to the history order.
 On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms.
@@
 (3) for the offline datasets before training.
-As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum value.
+As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum.
 Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms.
@@
 Process:
-1:  Initialize context $\hat{\bm{h}}_{0}=\{(\bm{x}_{0},y_{0},R_{0})\}$, where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding and $R_{0}=0$;
+1:  Initialize $\hat{\bm{h}}_{0}=\{(\bm{x}_{0},y_{0},R_{0})\}$, where $\bm{x}_{0}$ and $y_{0}$ are placeholders for padding and $R_{0}=0$;
 2:  for $t=1,2,\ldots,T$ do
@@
 6:     Relabel $R_{i}\leftarrow R_{i}+r$, for each $(\bm{x}_{i},y_{i},R_{i})$ in $\hat{\bm{h}}_{t-1}$;
-7:     $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},0)\}$;
+7:     $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},0)\}$
 8:  end for
@@
 A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$.
-Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and then decrease it as
-$R_{t}=R_{t-1}-(y^{*}-y_{t}).$ (5)
+Inspired by DT, a naive approach is to specify a desired performance as the initial RTG $R_{0}$, and decrease it as $R_{t}=R_{t-1}-(y^{*}-y_{t})$.
 However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
@@
 In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$.
-At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is first fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, and $\bm{x}_{t}$ is evaluated to obtain $y_{t}=f(\bm{x}_{t})$ in line 4.
-Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in the history $\hat{\bm{h}}_{t-1}$: calculating the instantaneous regret $r=y^{*}-y_{t}$ in line 5 and adding $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ in line 6, i.e.,
+At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, which is evaluated to obtain $y_{t}$ in line 4.
+Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ (line 6):
 $\forall 0\leq i<t,R_{i}\leftarrow R_{i}+(y^{*}-y_{t}).$ (5)
@@
 In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$.
-The above process is repeated until reaching the budget $T$.
-Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using zero as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over the future optimization history.
+This process is repeated until reaching the budget $T$.
+Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using 00 as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over future optimization history.
 3.3 Data Generation
@@
 Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection.
-For example, we can select some representatives of different types of BBO algorithms (e.g., BO and EA).
-Datasets $\mathcal{D}_{i,j}$ are then obtained by using each behavior algorithm to optimize each function with different random seeds.
-Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented using RTG tokens $R_{t}$ computed as in Eq.
+For example, we can select some representatives from different types of BBO algorithms, e.g., BO and EA.
+Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each function with different random seeds.
+Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$ computed as in Eq.
 (3).
-The resulting augmented histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for training the model $\mathcal{M}_{\bm{\theta}}$.
+The resulting histories $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T}$ compose the final datasets $\widehat{\mathcal{D}}_{i,j}$ for model training.
 Data Normalization.
@@
 We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively.
-For the function value $y$, we apply random scaling akin to previous works [61; 14].
+For the function value $y$, we apply random scaling akin to previous works [14, 64].
 That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$.
@@
 The model architecture and hyper-parameters are maintained consistently across these problems.
-We train our model using five distinct random seeds, ranging from 00 through $4$, and each trained model is also run five times independently during the execution phase.
+We train our model using five distinct random seeds, ranging from 00 to $4$, and each trained model is run five times independently during the execution phase.
 We will report the average performance and standard deviation.
-Details of the model hyper-parameters can be found in Appendix A.
-Our code is available at https://github.com/songlei00/RIBBO.
+Details of the model hyper-parameters are given in Appendix A.
+The codes are provided in the supplementary.
 4.1 Experimental Setup
@@
 For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties.
-According to the properties of the functions, they can be further divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
+According to the properties of the functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
 HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems.
@@
 For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation.
-Similar to [15; 56], we implement random translations and scalings to the search space to construct a distribution of functions.
+Similar to [15, 59], we implement random translations and scalings to the search space to construct a distribution of functions.
 For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors.
@@
 Data.
-Similar to OptFormer [14], we employ seven behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [49], Eagle Strategy [63], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively.
+Similar to OptFormer [14], we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [50], Eagle Strategy [66], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively.
 Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds.
@@
 4.2 Baselines
-As our proposed method RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with an offline dataset, including Behavior Cloning (BC) [7] and OptFormer [14].
+As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) [7] and OptFormer [14].
 Their hyper-parameters are set as same as that of our model for fairness.
-Note that the seven behavior algorithms used to generate datasets are also important baselines, which are included for comparison as well.
-Figure 2: Performance comparison among RIBBO, BC, BC Filter, OptFormer, and behavior algorithms on synthetic functions, HPO, and robot control problems.
-The $y$-axis is the normalized average objective value, and the length of vertical bars represents the standard deviation.
+Note that the seven behavior algorithms used to generate datasets are also important baselines, and included for comparison as well.
 BC uses the same transformer architecture as RIBBO.
@@
 To mitigate this issue, we have also trained the model by excluding these underperforming algorithms, denoted as BC Filter.
-OptFormer employs transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution.
+OptFormer employs a transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution.
 Its original implementation is built upon a text-based transformer with a large training scale.
-In this paper, we re-implemented a simplified version of OptFormer where we only maintain the algorithm type within the metadata, i.e., employing different initial states $\bm{x}_{0}$ and $y_{0}$ for different algorithms.
-More details about the training procedure and imitation capacity of the reimplementation can be found in Appendix C.
+In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm type within the metadata.
+Various initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms.
+They are obtained by indexing the algorithm type through an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer.
+Further training details can be found in Appendix C.
 4.3 Main Results
@@
 We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing zero as the immediate RTG to generate the optimal potential solutions.
-RIBBO does not perform well on SVM problem, which may be due to the problem’s low-dimensional nature (only three parameters) and its relative simplicity for optimization.
-Behavior algorithms can easily achieve good performance, while the complexity of RIBBO’s training and inference processes could instead result in the performance degradation.
-Regarding to rover problem where GP-EI performs the best, we collect less data from GP-EI than other behavior algorithms due to the high time cost.
+Figure 2: Performance comparison among RIBBO, BC, BC Filter, OptFormer, and behavior algorithms on synthetic functions, HPO, and robot control problems.
+The $y$-axis is the normalized average objective value, and the length of vertical bars represents the standard deviation.
+RIBBO does not perform well on the SVM problem, which may be due to the problem’s low-dimensional nature (only three parameters) and its relative simplicity for optimization.
+Behavior algorithms can achieve good performance easily, while the complexity of RIBBO’s training and inference processes could instead result in the performance degradation.
+For the rover problem where GP-EI performs the best, we collect less data from GP-EI than other behavior algorithms due to the high time cost.
 This may limit RIBBO’s capacity to leverage the high-quality data from GP-EI, given its small proportion relative to the data collected from other behavior algorithms.
 Despite this, RIBBO is still the runner-up, significantly surpassing the other behavior algorithms.
-(a) Visualization
-(b) Cross-dist generalization
-(c) Initial RTGs
-(d) RTG update strategy
-Figure 3: (a) Visualization of the contour lines of 2D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO.
-(b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties.
-Note that the legend shares with that of Figure 2.
-(c) Initial RTG $R_{0}$’s influence on performance.
-(d) RTG update strategy comparison between HRR and the naive strategy with different $R_{0}$.
 Compared with BC and BC Filter, RIBBO performs consistently better except on the SVM problem.
@@
 In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
+(a) Visualization
+(b) Cross-dist performance
+(c) Initial RTGs
+(d) RTG update strategy
+Figure 3: (a) Visualization of the contour lines of $2$D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO.
+(b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties.
+The legend shares with that of Figure 2.
+(c) Initial RTG $R_{0}$’s influence on performance.
+(d) RTG update strategy comparison between HRR and the naive strategy with various $R_{0}$.
 4.4 Ablation Studies
@@
 Influence of Initial RTG Token $R_{0}$.
-By incorporating RTG tokens into modeling, RIBBO should be able to attend to the RTGs and generate optimization trajectories that achieve the specified initial RTG token $R_{0}$.
-To validate this, we examine the performance of RIBBO when specifying different values of $R_{0}$, and the results are presented in Figure 3(c).
-Here, the RTG values, instead of the normalized objective values, of the trajectory are used as the $y$-axis.
-We can observe that the cumulated regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
+By incorporating RTG tokens, RIBBO can attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$.
+To validate this, we examine the performance of RIBBO with different values of $R_{0}$, and the results are presented in Figure 3(c).
+Here, the RTG values, instead of normalized objective values, are used as the $y$-axis.
+We can observe that the cumulative regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
 Effectiveness of HRR.
@@
 To assess the effectiveness of the proposed strategy HRR in Eq.
-(6), we compare it with the naive strategy in Eq.
-(5), which sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration.
+(5), we compare it with the naive strategy, that sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration.
 The results are shown in Figure 3(d).
 The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$.
-When $R_{0}=0$ i.e.
-the lower bound of regret, the model performs well in the initial stage.
+When $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially.
 However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values.
-Using $R_{0}=5$ compromises the initial performance, as the model may not choose the most aggressive solutions with a high $R_{0}$.
-However, higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in the late stage.
-The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG as 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq.
-(6) keeps the values meaningful and feasible.
-Other Ablation Studies.
-We also study the effect of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the length $\tau$ of sampled subsequences.
-Due to space limitation, we provide the results in Appendix E.
+Using $R_{0}=5$ compromises the initial performance, as the model may not select the most aggressive solutions with a high $R_{0}$.
+However, a higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage.
+The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq.
+(5) ensures the values remain meaningful and feasible.
+Further Studies.
+We also study the effects of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for $y$, the model size, and the sampled subsequence length $\tau$.
+Please see Appendix E.
+Generalization studies across the entire BBOB suite are shown in Appendix D, while the effects of transformations applied to the search space are detailed in Appendix F.
 5 Conclusion
@@
 For example, we only consider the model training over continuous search space with the same dimensionality, and it would be interesting to explore heteroscedastic search space with different types of variables.
-Potential Broader Impact
-This paper presents work whose goal is to advance the field of black-box optimization.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Acknowledgments and Disclosure of Funding
+Use unnumbered first level headings for the acknowledgments.
+All acknowledgments go at the end of the paper before the list of references.
+Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work).
+More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
+Do not include this section in the anonymized submission, only in the final paper.
+You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
 Appendix A Model Details
@@
 Details can be found in Table 1.
-For BC, BC Filter and OptFormer, the hyper-parameters are set as same as that of our model.
+For BC, BC Filter and OptFormer, the hyper-parameters are set as same as those of our model.
+The training takes about $20$ hours on $1$ GPU (Nvidia RTX $4090$).
 Table 1: List of hyper-parameter settings in RIBBO.
@@
 B.1 Benchmarks
-  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in continue domain.
+  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain.
 This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties.
@@
 We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5).
-We use the BBOB benchmark implementation in Open Source Vizier.
-The dimension is set to $10$ for all functions.
+We use the BBOB benchmark implementation in Open Source Vizier [^1] https://github.com/google/vizier, and the dimension is set to $10$ for all functions.
   • SVM and XGBoost.
-HPO-B [2] is the most commonly used HPO benchmark and grouped by search space id.
+HPO-B [2] is the most commonly used HPO benchmark and is grouped by search space id.
 Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost.
 Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset.
-For continue domain, it fits an XGBoost model as the objective function for each HPO problem.
+For the continuous domain, it fits an XGBoost model as the objective function for each HPO problem.
 These datasets for each search space id are divided into training and test datasets.
 We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively.
-SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most parameters in HPO-B.
-  • Rover Trajectory Planning [57; 16] is a trajectory optimization task to emulate a rover navigation task.
-The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, totally $60$ parameters to optimize.
-Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{x}_{s}\|_{1}+\|\bm{x}_{58,59}-\bm{x}_{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of a given trajectory, $\bm{x}_{s}$ and $\bm{x}_{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem.
+SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most in HPO-B.
+We use the official implementations [^2] https://github.com/releaunifreiburg/HPO-B.
+  • Rover Trajectory Planning [16, 60] is a trajectory optimization task designed to emulate a rover navigation task.
+The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, resulting in a total of $60$ parameters to optimize.
+Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{s}\|_{1}+\|\bm{x}_{58,59}-\bm{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of the given trajectory, $\bm{x}_{0,1}$ denotes the first and second dimensions of $\bm{x}$ (similarly for $\bm{x}_{58,59}$), $\bm{s}$ and $\bm{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem.
 This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions.
-To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ as well as scalings in $[0.9,1.1]$, similar to [56; 15].
+To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works [15, 59].
 The training and test tasks are randomly sampled from the distribution.
+We use the standard implementation for the rover problem [^3] https://github.com/uber-research/TuRBO.
 B.2 Bahavior Algorithms
-The datasets are generated with some representatives of heuristic search, EA, and BO as behavior algorithms.
+The datasets are generated with several representatives of heuristic search, EA, and BO as behavior algorithms.
 We use the implementation in Open Source Vizier [52] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES.
-For Hill Climbing and Regularized Evolution, we provide a simple reimplementation.
+For Hill Climbing and Regularized Evolution, we provide a simple re-implementation.
 We use the implementation in BoTorch [8] for GP-EI.
-The details of behavior algorithms are summarized as follows.
-  • Random Search uniformly at random selects a point in the domain at each iteration.
-  • Shuffled Grid Search discretizes the ranges of real parameters into $100$ equidistant points and then selects without replacement a random point from the grid at each iteration.
+The details of these behavior algorithms are summarized as follows.
+  • Random Search selects a point uniformly at random from the domain at each iteration.
+  • Shuffled Grid Search discretizes the ranges of real parameters into $100$ equidistant points and selects a random point from the grid without replacement at each iteration.
   • Hill Climbing.
-At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to $x_{next}$ and evaluated.
-If $f(x_{next})\textgreater f(x_{best})$, we update $x_{best}$ to $x_{next}$.
+At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to generate $x_{next}$, which is then evaluated.
+If $f(x_{next})\textgreater f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
   • Regularized Evolution [50] is an evolutionary algorithm with tournament selection and age-based replacement.
-We use a population size of $25$ and tournament size of $5$.
-At each iteration, we randomly select a tournament subset from the current population, and mutate the solution with maximum value.
-The mutation operation is to uniformly at random selects one of the parameters and mutates to a random value in the domain.
+We use a population size of $25$ and a tournament size of $5$.
+At each iteration, a tournament subset is randomly selected from the current population, and the solution with the maximum value is mutated.
+The mutation operation uniformly selects one of the parameters and mutates it to a random value within the domain.
   • Eagle Strategy [66] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies.
-Each firefly emits a light whose intensity corresponds to the objective value.
-At each iteration, for a firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker one in its vicinity, and the position is updated using the calculated weight.
+Each firefly emits light whose intensity corresponds to the objective value.
+At each iteration, for each firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker ones in its vicinity.
+The position is updated based on the calculated weight.
   • CMA-ES [26] is a popular evolutionary algorithm.
-At each iteration, the candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
+At each iteration, candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
   • GP-EI employs GP as the surrogate model and EI [30] as the acquisition function for BO.
 B.3 Data
-For training, a set of tasks are sampled from the task distribution and above behavior algorithms are used to collect data.
-Specifically, for BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times except GP-EI, which is limited to $100$ function samples due to high time cost.
+For training, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data.
+Specifically, for the BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times, except for GP-EI, which is limited to $100$ function samples due to its high time cost.
 For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost.
-All behavior algorithms employ $500$ random seeds in the execution except GP-EI, which is executed using $500$ random seeds.
-For rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times.
+All behavior algorithms employ $500$ random seeds in the execution.
+For the rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times.
 For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
@@
 OptFormer [14] is a general optimization framework based on transformer [57] and provides an interface to learn policy or function prediction.
-When provided with textual representations of historical data, OptFormer is capable of determining the next query point $\bm{x}_{t}$ as policy.
-Additionally, if the histories incorporate a possible query point $\bm{x}_{t}$, the framework is designed to predict the corresponding $y_{t}$, thereby serving as function prediction.
-We concentrate on the aspect of policy learning, attributable to its greater relevance to our work.
-The original implementation is built upon text-based transformer and uses private datasets for training.
-In this paper, we have re-implemented a simplified version of OptFormer, where we omit the textual tokenization process and only maintain the algorithm type within the metadata.
-Numerical inputs are fed into the model and we employ different initial states $\bm{x}_{0}$ and $y_{0}$ to represent different algorithms.
+When provided with textual representations of historical data, OptFormer can determine the next query point $\bm{x}_{t}$, acting as a policy.
+Additionally, if the context incorporate a possible query point $\bm{x}_{t}$, the framework is capable of predicting the corresponding $y_{t}$, thus serving as a prediction model.
+We focus on the aspect of policy learning, due to its greater relevance to our work.
+The original implementation is built upon a text-based transformer and uses private datasets for training.
+In this paper, we have re-implemented a simplified version of OptFormer, where we omit the textual tokenization process and only retain the algorithm type within the metadata.
+Numerical inputs are fed into the model and we use different initial states $\bm{x}_{0}$ and $y_{0}$ to distinguish among algorithms.
+The initial states are derived by indexing the algorithm type in an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer.
 The hyper-parameters are set as same as our method.
-To examine the algorithm imitation capability, we compare our re-implemented with the corresponding behavior algorithms.
+To examine the algorithm imitation capability, we compare our re-implementation with the corresponding behavior algorithms.
 The results are shown in Figure 4.
-For the sake of clarity in the visual representation, we only plot a subset of behavior algorithms, including Shuffled Grid Search, Hill Climbing, Regularized Evolution and Eagle Strategy.
+For clarity in the visual representation, we only plot a subset of the behavior algorithms, including Shuffled Grid Search, Hill Climbing, Regularized Evolution and Eagle Strategy.
 These behavior algorithms are plotted by solid lines, while their OptFormer counterparts are shown by dashed lines with the same color.
-Note that the figure is shown using the immediate function values as the $y$-axis to facilitate a comprehensive observation of the optimization.
+Note that the figure uses the immediate function values as the $y$-axis to facilitate a comprehensive observation of the optimization process.
 Figure 4: Comparison of the behavior algorithms with the OptFormer re-implementation.
-Appendix D Cross-Distribution Generalization
-We conduct experiments to examine the generalization of our method to unseen function distributions.
-We train the model on $4$ of $5$ chosen synthetic functions and test it with the remaining one.
+Appendix D Generalization
+To examine the generalization of our method to unseen function distributions, we conduct experiments to train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one.
 The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
-Figure 5: Cross-Distribution Generalization by training on $4$ synthetic functions and testing on the remaining one.
+Figure 5: Cross-distribution generalization by training on $4$ of $5$ chosen synthetic functions and testing on the remaining one.
+We also train the model across all training datasets from BBOB synthetic functions simultaneously, and evaluate on the same functions used in previous sections, specifically Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
+These functions serve as representatives of each category within the BBOB suite.
+The results are shown in Figure 6.
+RIBBO demonstrates relatively inferior performance compared to some behavior algorithms, which might be attributed to that the training dataset consists of excessive trajectories from diverse behavior algorithms across different functions, and complicates the process of algorithm identification solely through RTG tokens.
+Song et al.
+[53] had similar observation in Section 5.2 of their paper, indicating the significance of textual cues for $\bm{x}$ and the complete failure of the model when using anonymized inputs.
+However, RIBBO still surpasses both BC and BC Filter, emphasizing the effectiveness of RTG tokens.
+RIBBO employs RTG tokens to facilitate the identification of diverse algorithms across various functions and to generate sequences with specified regret.
+Further improvement could be achieved by integrating textual representations for $\bm{x}$ and metadata describing the problem, which may potentially improve the model’s ability to identify and generalize [54].
+This leaves as future work.
+Figure 6: Generalization by training across the entire BBOB suite.
 Appendix E Ablation Studies
-We provide further ablation studies to examine the influence of the important components and hyper-parameters of RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
-Token Aggregation aims to aggregate the information of $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them.
+We provide further ablation studies to examine the influence of the key components and hyper-parameters in RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
+Token Aggregation aims to aggregate the information from $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them.
 RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token.
-It is compared with two different token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave).
-The addition method integrates the values of each token to form a single token, while interleaving method addresses each token sequentially.
+This method is compared with two alternative token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave).
+The addition method integrates the values of each token into one, while the interleaving method addresses tokens sequentially.
 The results are shown in Figure 7(a).
-Concatenation method surpasses both the addition and interleaving methods, with the addition method exhibiting the least efficacy.
-The concatenation method employs a relatively straightforward technique to aggregate each token by direct concatenation, while the interleaving method adopts a more complex process to learn the interrelations of tokens.
-The inferior performance of the addition method is because that the summation will omit some details of the original tokens, which are necessary for the generation of the next query $\bm{x}_{t}$.
-Normalization Method is to balance the scales of function value $y$ across tasks.
-We compare the employed random normalization with the dataset normalization and no normalization.
-Dataset normalization scales the value $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values for $f_{i}$, while no normalization entails utilizing the original function values.
-The results are shown in Figure 6(b).
-Random normalization and dataset normalization have similar performance, while no normalization leads to the poor performance.
-The normalization method has a direct effect on the calculation of RTG tokens and thereby influences the generating with a desired regret.
-No normalization results in significant variances in the scale of $y$, posing a challenge for the generation process.
-Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating the training and inference.
-Nevertheless, random normalization can make a model exhibit invariance across various scales of $y$ as mentioned in [61; 14].
-Consequently, we recommend to use random normalization in practice.
+The concatenation method surpasses both the addition and interleaving methods, with the addition method showing the least efficacy.
+The concatenation method employs a relatively straightforward approach to aggregate tokens, while the interleaving method adopts a more complex process to learn the interrelations among tokens.
+The inferior performance of the addition method can be due to the summation operator, which potentially eliminates critical details from the original tokens.
+These details are essential for generating the subsequent query $\bm{x}_{t}$.
+Normalization Method is to balance the scales of the function value $y$ across tasks.
+We compare the employed random normalization with the dataset normalization and the absence of any normalization.
+Dataset normalization adjusts the value of $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for each task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values of $f_{i}$.
+No normalization uses the original function values directly.
+The results in Figure 7(b) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance.
+The choice of normalization method directly impacts the computation of RTG tokens and subsequently affects the generation of the desired regret.
+The lack of normalization leads to significant variations in the scale of $y$, which complicates the training process.
+Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating both training and inference.
+However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba and Grabocka [64] and Chen et al.
+[14].
+Therefore, we recommend using random normalization in practice.
 Model Size has an effect on the in-context learning ability.
-We compare the performance of the employed model size with a smaller one (8 layers, 4 attention heads and 128-dimensional embedding space) and a larger one (16 layers, 12 attention heads and 384-dimensional embedding space).
-The results are shown in Figure 6(c), indicating that the model’s size has small impact on its overall performance.
-A smaller model, with limited capacity, exhibits a diminished capability to the in-context learning ability, while a larger model requires more training data, potentially leading to a slight decrement in performance.
-Subsequence Length $\tau$ controls the length of context during both training and inference phases.
-Employing subsequence sampling can be viewed as a form of data augmentation that facilitates the training process.
-As shown in Figure 6(d), sampling subsequences instead of using the entire history as context, specifically when $\tau=T=150$, leads to an improvement of the performance.
-Causal transformer training and inference are quadratic in the context length, hence, a smaller $\tau$ can result in reduced computational complexity.
-However, an excessively shortened context may contain insufficient historical information, which could, in turn, adversely affect the performance.
+We assess the effects of different model sizes by comparing the performance of the currently employed model size with both a smaller and a larger model.
+The smaller model consists of $8$ layers, $4$ attention heads, and $128$-dimensional embedding space, while the larger model has $16$ layers, $12$ attention heads, and $384$-dimensional embedding space.
+The results are shown in Figure 7(c), indicating that the model size has a minimal impact on overall performance.
+Specifically, the smaller model, due to its limited capacity, shows a reduced performance, while the larger one, potentially more powerful, requires more training data, which can lead to a slight decrease in performance due to overfitting or inefficiency in learning from a limited dataset.
+This analysis highlights the trade-offs involved in selecting the appropriate model size for optimal performance.
+Subsequence Length $\tau$ controls the context length during both the training and inference phases.
+The process of subsequence sampling acts as a form of data augmentation, enhancing training efficiency.
+As shown in Figure 7(d), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance.
+The computational complexity for causal transformer training and inference scales quadratically with the context length.
+Therefore, utilizing a shorter $\tau$ can significantly reduce computational demands.
+However, it is important to note that a shorter $\tau$ might not capture sufficient historical data, potentially degrading the performance due to the insufficient contextual information.
+This highlights the trade-offs between computational complexity and performance.
 (a) Token aggregation
@@
 (d) Subsequence length $\tau$
-Figure 6: Ablation studies of token aggregation, normalization method, model size, and the subsequence length $\tau$.
+Figure 7: Ablation studies of token aggregation, normalization method, model size, and subsequence length $\tau$.
+Appendix F Visualization of Branin Function
+In several of our experiments, such as BBOB, rover problems, and the visualization analysis of the Branin function, we implement a series of transformations to the search space.
+This is designed to generate a distribution of functions with similar properties.
+A set of functions is sampled serving as the training and test tasks from the distribution.
+For the Branin function, random translations and scalings are applied to form the distribution.
+In this section, we present visualizations of the contour lines of the $2$D Branin functions, sampled from the distribution to demonstrate the effects of the applied transformations.
+To this end, $4$ distinct random seeds are used to draw samples from the distribution.
+The visualizations are shown in Figure 8.
+Note that the two parameters of the Branin function have been scaled to the range of $[-1,1]$ for the clarity of visual representations.
+Variations are observed in both the location of the optimum and the scales of the objective values.
+For instance, the optimum of the first subfigure locates more to the right compared to the second one, and the contour lines in the former are much more dispersed than those in the latter.
+The value scale for the first subfigure ranges from $-8$ to $1$, while in the second, it ranges from $-3.5$ to $1$.
+By applying these transformations, we can generate a function distribution that retains similar properties, enabling the sampling of training and test tasks for our model.
+For other functions, such as BBOB suite, besides simple random translations and scalings, more complex transformations such as non-linear Tosz and Tasy transformations are applied, leading to a more intricate landscape.
+However, due to the high dimensionality of these functions, direct visualization is impractical, so we present only the visualization of the $2$D Branin function.
+Figure 8: Contour line visualization for samples drawn from the $2$D Branin function distribution using $4$ distinct random seeds.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The abstract and introduction clearly state the claims made, and the experimental results reflect the effectiveness of our method.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: See the last paragraph of the paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: See Appendix A and B. The codes are provided in the supplementary.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The codes are provided in the supplementary and the scripts to reproduce experimental results are provided.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The experimental setting is presented in Section 4, and the full details are provided in Appendix A and B. The codes are provided in the supplementary.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We show error bars by the length of vertical bars in the figures.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: See Appendix A.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have reviewed the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: There is no societal impact of the work performed.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: See Appendix B.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: The paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
 1 Introduction
-Black-Box Optimization (BBO) [1, 4] refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available.
+Black-Box Optimization (BBO) Alarie et al.
+(2021); Audet & Hare (2017) refers to optimizing objective functions where neither analytic expressions nor derivatives of the objective are available.
 To solve BBO problems, we can only access the results of objective evaluation, which usually also incurs a high computational cost.
-Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery [42, 55], material design [21, 23], robot control [10, 11], and optimal experimental design [24, 44], just to name a few.
-To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) [20, 51] and Evolutionary Algorithms (EA) [5, 68].
+Many fundamental problems in science and engineering involve optimization of expensive BBO functions, such as drug discovery Negoescu et al.
+(2011); Terayama et al.
+(2021), material design Frazier & Wang (2016); Gómez-Bombarelli et al.
+(2018), robot control Calandra et al.
+(2016); Chatzilygeroudis et al.
+(2019), and optimal experimental design Greenhill et al.
+(2020); Nguyen et al.
+(2023), just to name a few.
+To date, a lot of BBO algorithms have been developed, among which the most prominent ones are Bayesian Optimization (BO) Frazier (2018); Shahriari et al.
+(2016) and Evolutionary Algorithms (EA) Back (1996); Zhou et al.
+(2019).
 Despite the advancements, these algorithms typically solve BBO problems from scratch and rely on expert-derived heuristics.
-Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems [3, 6].
-Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data [2, 18].
+Consequently, they are often hindered by slow convergence rates, and unable to leverage the inherent structures within the optimization problems Astudillo & Frazier (2021); Bai et al.
+(2023).
+Recently, there has been a growing interest in meta-learning a particular component of the algorithms with previously collected data Arango et al.
+(2021); Feurer et al.
+(2021).
 Learning the component not only alleviates the need for the laborious design process of the domain experts, but also specifies the component with domain data to facilitate subsequent optimization.
-For example, some components in BO are proposed to be learned from data, including the surrogate model [41, 46, 61, 64], acquisition function [29, 59], initialization strategy [17, 47], and search space [37, 45]; some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm [34] or the update rules for evolution strategy [35].
+For example, some components in BO are proposed to be learned from data, including the surrogate model Müller et al.
+(2023); Perrone et al.
+(2018); Wang et al.
+(2021); Wistuba & Grabocka (2021), acquisition function Hsieh et al.
+(2021); Volpp et al.
+(2020), initialization strategy Feurer et al.
+(2015); Poloczek et al.
+(2016), and search space Li et al.
+(2022); Perrone & Shen (2019); some core evolutionary operations in EA have also been considered, e.g., learning the selection and mutation rate adaptation in genetic algorithm Lange et al.
+(2023a) or the update rules for evolution strategy Lange et al.
+(2023b); the configuration of the algorithm can also be learned and dynamically adjusted throughout the optimization process Biedenkapp et al.
+(2020); Adriaensen et al.
+(2022).
 There have also been some attempts to learn an entire algorithm in an End-to-End (E2E) fashion, which requires almost no expert knowledge at all and provides the most flexibility across a broad range of BBO problems.
-However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) [13, 56] or online sampling from the objective function (often very expensive) [39]Chen et al.
-[14] proposed the OptFormer method to imitate the behavior algorithms separately during training, requiring a user to manually specify which algorithm to execute during testing, which still poses a challenge.
+However, existing practices require additional knowledge regarding the objective function during the training stage, e.g., the gradient information (often impractical for BBO) Chen et al.
+(2017); TV et al.
+(2019) or online sampling from the objective function (often very expensive) Maraval et al.
+(2023)Chen et al.
+(2022) proposed the OptFormer method to imitate the behavior algorithms separately during training, presenting a challenge for the user to manually specify which algorithm to execute during testing.
 Thus, these methods are less ideal for practical scenarios where offline datasets are often available beforehand and a suitable algorithm for the given task has to be identified automatically without the involvement of domain experts.
@@
 The sequence model is fed with previous query points and their function values, and trained to predict the distribution over the next query point.
-During testing, the sequence model itself serves as a BBO algorithm by generating the query points auto-regressively.
+During testing, the sequence model itself serves as a BBO algorithm by generating the next query points auto-regressively.
 Apart from this, RIBBO augments the optimization histories with regret-to-go (RTG) tokens, which are calculated by summing up the regrets over the future part of the histories, representing the future performance of an algorithm.
@@
 Note that the best behavior algorithm depends on the problem at hand, and RIBBO can perform even better on some problems.
-Compared to the most related method OptFormer [14], RIBBO also has clear advantage.
+Compared to the most related method OptFormer Chen et al.
+(2022), RIBBO also has clear advantage.
 In addition, we perform a series of experiments to analyze the influence of important components of RIBBO.
@@
 The goal of BBO is to find an optimal solution $\bm{x}^{*}\in\mathop{\arg\max}_{\bm{x}\in\mathcal{X}}f(\bm{x})$, with the only permission of querying the objective function value.
-Several classes of BBO algorithms have been proposed, e.g., BO [20, 51] and EA [5, 68].
-The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) [49], and an acquisition function [63], which are used to model $f$ and decide the next query point, respectively.
+Several classes of BBO algorithms have been proposed, e.g., BO Frazier (2018); Shahriari et al.
+(2016) and EA Back (1996); Zhou et al.
+(2019).
+The basic framework of BO contains two critical components: a surrogate model, typically formalized as Gaussian Process (GP) Rasmussen & Williams (2006), and an acquisition function Wilson et al.
+(2018), which are used to model $f$ and decide the next query point, respectively.
 EA is a class of heuristic optimization algorithms inspired by natural evolution.
@@
 2.2 Meta-Learning in Black-Box Optimization
-Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior on a wide range of problems, which is tedious and time-consuming.
-One solution is meta-learning [28, 58], which aims to exploit knowledge to improve the performance of learning algorithms given data from a set of tasks.
-By parameterizing a component of BBO algorithms or even an entire algorithm that is traditionally manually designed, we can learn from historical data to incorporate the knowledge into the optimization, which may bring speedup.
+Hand-crafted BBO algorithms usually require an expert to analyze the algorithms’ behavior across a wide range of problems, a process that is both tedious and time-consuming.
+One solution is meta-learning Vilalta & Drissi (2002); Hospedales et al.
+(2021), which aims to exploit knowledge to improve the performance of learning algorithms given data from a collection of tasks.
+By parameterizing a component of BBO algorithms or even an entire BBO algorithm that is traditionally manually designed, we can utilize historical data to incorporate the domain knowledge into the optimization, which may bring speedup.
 Meta-learning particular components has been studied with different BBO algorithms.
-Meta-learning in BO can be divided into four main categories according to “what to transfer” [6], including the design of the surrogate model, acquisition function, initialization strategy, and search space.
+Meta-learning in BO can be divided into four main categories according to “what to transfer” Bai et al.
+(2023), including the design of the surrogate model, acquisition function, initialization strategy, and search space.
 For surrogate model design, Wang et al.
-[61] and Wistuba and Grabocka [64] parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al.
-[46] and Müller et al.
-[41] substituted GP with Bayesian linear regression or neural process [22, 40].
-For acquisition function design, MetaBO [59] uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF [29] employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning [19].
-The remaining two categories focus on exploiting the previous good solutions to warm start the optimization [17, 47] or shrink the search space [37, 45].
-Meta-learning in EA usually focuses on specific evolutionary operations.
+(2021) and Wistuba & Grabocka (2021) parameterized the mean or kernel function of the GP model with Multi-Layer Perceptron (MLP), while Perrone et al.
+(2018) and Müller et al.
+(2023) substituted GP with Bayesian linear regression or neural process Garnelo et al.
+(2018); Müller et al.
+(2022).
+For acquisition function design, MetaBO Volpp et al.
+(2020) uses Reinforcement Learning (RL) to meta-train an acquisition function on a set of related tasks, and FSAF Hsieh et al.
+(2021) employs a Bayesian variant of deep Q-network as a surrogate differentiable acquisition function trained by model-agnostic meta-learning Finn et al.
+(2017).
+The remaining two categories focus on exploiting the previous good solutions to warm start the optimization Feurer et al.
+(2015); Poloczek et al.
+(2016) or shrink the search space Perrone & Shen (2019); Li et al.
+(2022).
+Meta-learning in EA usually focuses on learning specific evolutionary operations.
 For example, Lang et al.
-substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules [34], and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules [35].
+substituted core genetic operators, i.e., selection and mutation rate adaptation, with dot-product attention modules Lange et al.
+(2023a), and meta-learned a self-attention-based architecture to discover effective and order-invariant update rules Lange et al.
+(2023b).
+ALDes Zhao et al.
+(2024) introduces an auto-regressive learning-based approach to sequentially generate components of meta-heuristic algorithms.
+Beyond that, Dynamic Algorithm Configuration (DAC) Biedenkapp et al.
+(2020); Adriaensen et al.
+(2022) concentrates on learning the configurations of algorithms, employing RL to dynamically adjust the configurations during the optimization process.
 Meta-learning entire algorithms has also been explored to obtain more flexible models.
-Early works [13, 56] use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations.
+Early works Chen et al.
+(2017); TV et al.
+(2019) use Recurrent Neural Network (RNN) to meta-learn a BBO algorithm by optimizing the summed objective functions of some iterations.
 RNN uses its memory state to store information about history and outputs the next query point.
-This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO.
-OptFormer [14] uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems.
-It can simultaneously imitate at least $7$ different BBO algorithms on a broad range of problems, which, however, requires a user to specify an algorithm manually when applying it to solve problems.
-Thus, how to select suitable algorithms automatically is still a challenge.
-Neural Acquisition Processes (NAP) [39] uses transformer to meta-learn the surrogate model and acquisition function of BO jointly.
+This work assumes access to gradient information during the training phase, which is, however, usually impractical in BBO problems.
+OptFormer Chen et al.
+(2022) uses a text-based transformer framework to learn an algorithm, providing a universal E2E interface for BBO problems.
+It is trained to imitate different BBO algorithms across a broad range of problems, which, however, presents a challenge for the user to manually specify an algorithm for inference.
+Neural Acquisition Processes (NAP) Maraval et al.
+(2023) uses transformer to meta-learn the surrogate model and acquisition function of BO jointly.
 Due to the lack of labeled acquisition data, NAP uses an online RL algorithm with a supervised auxiliary loss for training, which requires online sampling from the expensive objective function and lacks efficiency.
-Black-box Optimization NETworks (BONET) [33] employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing.
+Black-box Optimization NETworks (BONET) Krishnamoorthy et al.
+(2023) employ a transformer model to fit regret-augmented trajectories in an offline BBO scenario, where the training and testing data are from the same objective function, and a prefix sequence is required to warm up the optimization before testing.
 Compared to the above state-of-the-art E2E methods, we consider the meta-BBO setting, where the training datasets consist of diverse algorithms across different functions.
-Our approach offers the advantage of automatically identifying and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase, and utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
+Our approach offers the advantage of automatically identifying with RTG tokens and deploying the best-performing algorithm without requiring the user to pre-specify which algorithm to use or to provide a prefix sequence during the testing phase.
+It utilizes a supervised learning loss for training on a fixed offline dataset without the need for further interaction with the objective function.
 2.3 Decision Transformer
-Transformer has emerged as a powerful architecture for sequence modeling tasks [32, 62, 65].
-A basic building block behind transformer is the self-attention mechanism [57], which captures the correlation between tokens of any pair of timesteps.
-As the scale of data and model increases, transformer has demonstrated the in-context learning ability [9], which refers to the capability of the model to infer the tasks at hand based on the input contexts.
-Decision Transformer (DT) [12] abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions.
+Transformer has emerged as a powerful architecture for sequence modeling tasks Khan et al.
+(2022); Wen et al.
+(2022); Wolf et al.
+(2020).
+A basic building block behind transformer is the self-attention mechanism Vaswani et al.
+(2017), which captures the correlation between tokens of any pair of timesteps.
+As the scale of data and model increases, transformer has demonstrated the in-context learning ability Brown et al.
+(2020), which refers to the capability of the model to infer the tasks at hand based on the input contexts.
+Decision Transformer (DT) Chen et al.
+(2021) abstracts RL as a sequence modeling problem, and introduces return-to-go tokens, representing the cumulative rewards over future interactions.
 Conditioning on return-to-go tokens enables DT to correlate the trajectories with their corresponding returns and generate future actions to achieve a user-specified return.
@@
 This section presents Reinforced In-context Black-Box Optimization (RIBBO), which learns an enhanced BBO algorithm in an E2E fashion, as illustrated in Figure 1.
-We follow the task-distribution assumption, which is commonly adopted in meta-learning settings [19, 28, 67].
+We follow the task-distribution assumption, which is commonly adopted in meta-learning settings Finn et al.
+(2017); Hospedales et al.
+(2021); Zhou et al.
+(2023).
 Our goal is to learn a generalizable model $\mathcal{M}$ capable of solving a wide range of BBO tasks, each associated with a BBO objective function $f$ sampled from the task distribution $P(\mathcal{F})$, where $\mathcal{F}$ denotes the function space.
@@
 Given the current history $\bm{h}_{t-1}$ at iteration $t$, a BBO algorithm selects the next query point $\bm{x}_{t}$, observes the function value $y_{t}=f_{i}(\bm{x}_{t})$, and updates the history $\bm{h}_{t}=\bm{h}_{t-1}\cup\{(\bm{x}_{t},y_{t})\}$.
-Similar to the previous work [13], we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$.
-The optimization histories in offline datasets provide a natural supervision for learning.
-Given a set of histories $\{\bm{h}_{T}\}$ generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$, we employ a causal transformer model $\mathcal{M}_{\bm{\theta}}$, and expect it to imitate $\mathcal{A}$ and produce similar optimization history on $f$.
+Similar to the previous work Chen et al.
+(2017), we take this framework as a starting point and treat the learning of a universal BBO algorithm as learning a model $\mathcal{M}_{\bm{\theta}}$, which takes the preceding history $\bm{h}_{t-1}$ as input and outputs a distribution of the next query point $\bm{x}_{t}$.
+The optimization histories in offline datasets provide a natural supervision for the learning process.
+Suppose we have a set of histories $\{\bm{h}_{T}\}$, generated by a single behavior algorithm $\mathcal{A}$ on a single task $f$.
+By employing a causal transformer model $\mathcal{M}_{\bm{\theta}}$, we expect $\mathcal{M}_{\bm{\theta}}$ to imitate $\mathcal{A}$ and produce similar optimization history on $f$.
 In practice, we usually have datasets containing histories from multiple behavior algorithms $\{\mathcal{A}_{j}\}_{j=1}^{K}$ on multiple tasks $\{f_{i}\}_{i=1}^{N}$.
@@
 Nevertheless, naively imitating the offline datasets hinders the model since some inferior behavior algorithms may severely degenerate the model’s performance.
-Inspired by DT [12], we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
+Inspired by DT Chen et al.
+(2021), we propose to augment the optimization history with Regret-To-Go (RTG) tokens $R_{t}$, defined as the sum of instantaneous regrets over the future history:
 $\hat{\bm{h}}_{T}=\{(\bm{x}_{t},y_{t},R_{t})\}_{t=0}^{T},\;R_{t}=\sum\nolimits_{t^{\prime}=t+1}^{T}(y^{*}-y_{t^{\prime}}),$ (3)
@@
 The resulting method RIBBO has implicitly utilized the in-context learning capacity of transformer to guide the optimization with previous histories and the desired future regret as context.
-The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases [31].
+The in-context learning capacity of inferring the tasks at hand based on the input contexts has been observed as the scale of data and model increases Kaplan et al.
+(2020).
 It has been explored to infer general functional relationships as supervised learning or RL algorithms.
-For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context [25, 27, 38]Laskin et al.
-[36] learned RL algorithms using causal transformers.
+For example, the model is expected to predict accurately on the query input $\bm{x}_{t}$ by feeding the training dataset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{t-1}$ as the context Guo et al.
+(2023); Hollmann et al.
+(2023); Li et al.
+(2023)Laskin et al.
+(2023) learned RL algorithms using causal transformers.
 Here, we use it for BBO.
@@
 Model Architecture.
-For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture [48], which comprises a stack of causal attention blocks, each containing an attention component and a feed-forward network.
+For the formalization of the model $\mathcal{M}_{\bm{\theta}}$, we adopt the commonly used GPT architecture Radford et al.
+(2018), which comprises a stack of causal attention blocks.
+Each block is composed of an attention mechanism and a feed-forward network.
 We aggregate each triplet $(\bm{x}_{i},y_{i},R_{i})$ using a two-layer MLP network.
 The output of $\mathcal{M}_{\bm{\theta}}$ is a diagonal Gaussian distribution of the next query point.
-Note that previous works that adopt the sequence model as a surrogate model [40, 41, 43] typically remove the positional encoding because the surrogate model should be invariant to the history order.
+Note that previous works that adopt the sequence model as surrogate models Müller et al.
+(2022; 2023); Nguyen & Grover (2022) typically remove the positional encoding because the surrogate model should be invariant to the history order.
 On the contrary, our implementation preserves the positional encoding, naturally following the behavior of certain algorithms (e.g., BO or EA) and making it easier to learn from algorithms.
-Additionally, the positional encoding can keep the monotonically decreasing order of RTG tokens.
-More architecture details can be found in Appendix A.
+Additionally, the positional encoding can help maintain the monotonically decreasing order of RTG tokens.
+More details about the architecture can be found in Appendix A.
 Model Training.
-RTG tokens are computed as outlined in Eq.
+RTG tokens are calculated as outlined in Eq.
 (3) for the offline datasets before training.
-As the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum.
+Since the calculation of regret requires the optimum value of task $i$, we use the best-observed value $y_{\rm max}^{i}$ as a proxy for the optimum.
 Let $\{\widehat{\mathcal{D}}_{i,j}\}_{i\in[N],j\in[K]}$ denote the RTG augmented datasets with $N$ tasks and $K$ algorithms.
-We sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets and train $\mathcal{M}_{\bm{\theta}}$ to minimize the RIBBO loss in Eq.
+During training, we sample a minibatch of consecutive subsequences of length $\tau<T$ uniformly from the augmented datasets.
+The training objective is to minimize the RIBBO loss in Eq.
 (4).
@@
 Model Inference.
-The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during the inference process, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$.
+The model $\mathcal{M}_{\bm{\theta}}$ generates the query points $\bm{x}_{t}$ auto-regressively during inference, which involves iteratively selecting a new query point $\bm{x}_{t}$ based on the current augmented history $\hat{\bm{h}}_{t-1}$, evaluating the query as $y_{t}=f(\bm{x}_{t})$, and updating the history by $\hat{\bm{h}}_{t}=\hat{\bm{h}}_{t-1}\cup\{(\bm{x}_{t},y_{t},R_{t})\}$.
 A critical aspect of this process is how to specify the value of RTG (i.e., $R_{t}$) at every iteration $t$.
@@
 However, this strategy has the risk of producing out-of-distribution RTGs, since the values can fall below 00 due to an improperly selected $R_{0}$.
-Given the fact that RTGs are lower bounded by 00 and a 00 value implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 0, and use a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sampling.
+Given the fact that RTGs are lower bounded by 00 and a value of 00 implies a good BBO algorithm with low regret, we propose to set the immediate RTG as 00.
+Furthermore, we introduce a strategy called Hindsight Regret Relabelling (HRR) to update previous RTGs based on the current sample evaluations.
 The inference procedure with HRR is detailed in Algorithm 1.
 In line 1, the history $\hat{\bm{h}}_{0}$ is initialized with padding placeholders $\bm{x}_{0},y_{0}$ and RTG $R_{0}=0$.
-At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, which is evaluated to obtain $y_{t}$ in line 4.
-Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token in $\hat{\bm{h}}_{t-1}$ (line 6):
+At iteration $t$ (i.e., lines 3–7), the model $\mathcal{M}_{\bm{\theta}}$ is fed with the augmented history $\hat{\bm{h}}_{t-1}$ to generate the next query point $\bm{x}_{t}$ in line 3, followed by the evaluation procedure to obtain $y_{t}$ in line 4.
+Then, the immediate RTG $R_{t}$ is set to 00, and we employ HRR to update previous RTG tokens in $\hat{\bm{h}}_{t-1}$, i.e., calculate the instantaneous regret $r=y^{*}-y_{t}$ (line 5) and add $r$ to every RTG token within $\hat{\bm{h}}_{t-1}$ (line 6):
 $\forall 0\leq i<t,R_{i}\leftarrow R_{i}+(y^{*}-y_{t}).$ (5)
@@
 In line 7, the history $\hat{\bm{h}}_{t}$ is updated by expanding $\hat{\bm{h}}_{t-1}$ with $\{(\bm{x}_{t},y_{t},0)\}$, i.e., the current sampling and its immediate RTG $R_{t}=0$.
-This process is repeated until reaching the budget $T$.
+The above process is repeated until reaching the budget $T$.
 Thus, we can find that HRR not only exploits the full potential of $\mathcal{M}_{\bm{\theta}}$ through using 00 as the immediate RTG and thereby demands the model to generate the most advantageous decisions, but also preserves the calculation of RTG tokens following the same way as the training data, i.e., representing the cumulative regret over future optimization history.
@@
 Data Collection.
-Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can select a diverse set of algorithms for data collection.
+Given a set of tasks $\{f_{i}\}_{i=1}^{N}$ sampled from the task distribution $P(\mathcal{F})$, we can employ a diverse set of behavior algorithms for data collection.
 For example, we can select some representatives from different types of BBO algorithms, e.g., BO and EA.
-Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each function with different random seeds.
-Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$ computed as in Eq.
+Datasets $\mathcal{D}_{i,j}$ are obtained by using each behavior algorithm to optimize each task with different random seeds.
+Each optimization history $\bm{h}_{T}=\{(\bm{x}_{t},y_{t})\}_{t=1}^{T}$ in $\mathcal{D}_{i,j}$ is then augmented with RTG tokens $R_{t}$, which is computed as in Eq.
 (3).
@@
 We normalize the point $\bm{x}$ by $(\bm{x}-\bm{x}_{\rm min})/(\bm{x}_{\rm max}-\bm{x}_{\rm min})$, with $\bm{x}_{\rm max}$ and $\bm{x}_{\rm min}$ being the upper and lower bounds of the search space, respectively.
-For the function value $y$, we apply random scaling akin to previous works [14, 64].
-That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$.
+For the function value $y$, we apply random scaling akin to previous works Wistuba & Grabocka (2021); Chen et al.
+(2022).
+That is, when sampling a history $\bm{h}_{\tau}$ from the datasets $\mathcal{D}_{i,j}$, we randomly sample the lower bound $l\sim\mathcal{U}(y^{i}_{\rm min}-\frac{s}{2},y^{i}_{\rm min}+\frac{s}{2})$ and the upper bound $u\sim\mathcal{U}(y^{i}_{\rm max}-\frac{s}{2},y^{i}_{\rm max}+\frac{s}{2})$, where $\mathcal{U}$ stands for uniform distribution, $y^{i}_{\rm min},y^{i}_{\rm max}$ denote the observed minimum and maximum values for $f_{i}$, and $s=y^{i}_{\rm max}-y^{i}_{\rm min}$; the values $y_{t}$ in $\bm{h}_{\tau}$ are then normalized by $(y_{t}-l)/(u-l)$ for training.
 The RTG tokens are calculated accordingly with the normalized values.
 The random normalization can make a model exhibit invariance across various scales of $y$.
+For inference, the average values of the best-observed and worst-observed values across the training tasks are used to normalize $y$.
 4 Experiments
@@
 Details of the model hyper-parameters are given in Appendix A.
-The codes are provided in the supplementary.
+Our code is available at https://github.com/songlei00/RIBBO.
 4.1 Experimental Setup
 Benchmarks.
-We use BBO Benchmarks BBOB [15], HPO-B [2], and rover trajectory planning tasks [60].
+We use BBO Benchmarks BBOB Elhara et al.
+(2019), HPO-B Arango et al.
+(2021), and rover trajectory planning task Wang et al.
+(2018).
 The BBOB suite, a comprehensive and widely used benchmark in the continuous domain, consists of $24$ synthetic functions.
-For each function, a series of transformations are implemented on the search space to obtain a distribution of functions with similar properties.
-According to the properties of the functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
-HPO-B is the most commonly used HPO benchmark and consists of a series of HPO problems.
-It fits an XGBoost model as the objective function for each HPO problem, and we conduct experiments on two widely used machine models, SVM and XGBoost.
-For robot control optimization, we perform experiments on rover trajectory planning, which is a trajectory optimization task to emulate rover navigation.
-Similar to [15, 59], we implement random translations and scalings to the search space to construct a distribution of functions.
-For BBOB and rover problems, we sample a set of functions as training and test tasks, while for HPO-B, we use the meta-training/test splits provided by the authors.
+For each function, a series of linear and non-linear transformations are implemented on the search space to obtain a distribution of functions with similar properties.
+According to the properties of these functions, they can be divided into $5$ categories, and we select one from each category due to resource constraints, including Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
+HPO-B is a commonly used HPO benchmark and consists of a series of HPO problems.
+Each problem is to optimize a machine learning model across various datasets, and an XGBoost model is provided as the objective function for evaluation in a continuous space.
+We conduct experiments on two widely used models, SVM and XGBoost, in the continue domain.
+For robot control optimization, we perform experiments on rover trajectory planning task, which is a trajectory optimization problem to emulate rover navigation.
+Similar to Elhara et al.
+(2019); Volpp et al.
+(2020), we implement random translations and scalings to the search space to construct a distribution of functions.
+For BBOB and rover problems, we sample a set of functions from the task distribution as training and test tasks, while for HPO-B, we use the meta-training/test task splits provided by the authors.
 Detailed explanations of the benchmarks can be found in Appendix B.1.
 Data.
-Similar to OptFormer [14], we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution [50], Eagle Strategy [66], CMA-ES [26], and GP-EI [8], which are representatives of heuristic search, EA, and BO, respectively.
-Datasets are generated by using each behavior algorithm to optimize each training function with different random seeds.
-More details about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
+Similar to OptFormer Chen et al.
+(2022), we employ $7$ behavior algorithms, i.e., Random Search, Shuffled Grid Search, Hill Climbing, Regularized Evolution Real et al.
+(2019), Eagle Strategy Yang & Deb (2010), CMA-ES Hansen (2016), and GP-EI Balandat et al.
+(2020), which are representatives of heuristic search, EA, and BO, respectively.
+Datasets are generated by employing each behavior algorithm to optimize various training functions sampled from the task distribution, using different random seeds.
+For inference, new test functions are sampled from the same task distribution to serve as the test set.
+Specifically, for the HPO-B problem, the meta-training/test splits have been predefined by the authors and we adhere to this standard setup.
+Additional information about the behavior algorithms and datasets can be found in Appendix B.2 and B.3.
 4.2 Baselines
-As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) [7] and OptFormer [14].
+As RIBBO is an in-context E2E model, the most related baselines are those also training an E2E model with offline datasets, including Behavior Cloning (BC) Bain & Sammut (1995) and OptFormer Chen et al.
+(2022).
 Their hyper-parameters are set as same as that of our model for fairness.
@@
 To mitigate this issue, we have also trained the model by excluding these underperforming algorithms, denoted as BC Filter.
-OptFormer employs a transformer to imitate the behaviors of a set of algorithms and needs to manually specify the imitated algorithm during execution.
+OptFormer employs a transformer to imitate the behaviors of a set of algorithms and an algorithm identifier usually needs to be specified manually during inference for superior performance.
 Its original implementation is built upon a text-based transformer with a large training scale.
-In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm type within the metadata.
-Various initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms.
-They are obtained by indexing the algorithm type through an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer.
-Further training details can be found in Appendix C.
+In this paper, we re-implement a simplified version of OptFormer where we only retain the algorithm identifier within the metadata.
+The initial states, denoted as $\bm{x}_{0}$ and $y_{0}$, are used to distinguish between algorithms.
+They are obtained by indexing the algorithm type through an embedding layer, thereby aligning the initial states with the specific imitated algorithm.
+This enables the identification of distinct behavior algorithms within the simplified OptFormer.
+Further details about the re-implementation can be found in Appendix C.
 4.3 Main Results
@@
 It is because RIBBO is capable of using RTG tokens to identify algorithms and reinforce the performance on top of the behavior algorithms, which will be clearly shown later.
-We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing zero as the immediate RTG to generate the optimal potential solutions.
+We can also observe that RIBBO performs extremely well in the early stage, which draws the advantage from the HRR strategy, i.e., employing 00 as the immediate RTG to generate the optimal potential solutions.
 Figure 2: Performance comparison among RIBBO, BC, BC Filter, OptFormer, and behavior algorithms on synthetic functions, HPO, and robot control problems.
@@
 Compared with BC and BC Filter, RIBBO performs consistently better except on the SVM problem.
-BC tends to imitate the average behavior of various algorithms, and its poor performance is due to the aggregation of data from behavior algorithms with inferior performance.
+BC tends to imitate the average behavior of various algorithms, and its poor performance is due to the aggregation of behavior algorithms with inferior performance.
 BC Filter is generally better than BC, because the data from the two underperforming behavior algorithms, i.e., Random Search and Shuffled Grid Search, are manually excluded for the training of BC Filter.
@@
 Additionally, RIBBO achieves the exploration and exploitation trade-off capability using its knowledge about the task obtained during training, thus generating superior solutions over the ones in the training dataset.
-Cross-Distribution Generalization.
-We further conduct experiments to examine the generalization of RIBBO to unseen function distributions.
-For this purpose, we train the model on $4$ of the $5$ chosen synthetic functions and test it with the remaining one.
-Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and we sample a set of functions from each distribution as introduced before.
-Greiwank Rosenbrock is used for testing.
-The results in Figure 3(b) suggest that RIBBO has a strong generalizing ability to unseen function distributions.
-More results on cross-distribution generalization are deferred to Appendix D.
-The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in [13].
+Generalization.
+We also conduct experiments to examine the generalization capabilities of RIBBO.
+For this purpose, we train the model on all $24$ BBOB synthetic functions simultaneously with the results shown in Figure 3(b).
+To aggregate results across functions with different output scaling, we normalize all the functions adhering to previous literature Turner et al.
+(2020); Arango et al.
+(2021); Chen et al.
+(2022).
+The results suggest that RIBBO demonstrates strong generalizing to a variety of functions with different properties.
+Further experiments are conducted to examine the cross-distribution generalization to unseen function distributions.
+The model is trained on $4$ of the $5$ chosen synthetic functions and tested with the remaining one.
+Note that each function (i.e., Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge) here actually represents a distribution of functions with similar properties, and a set of functions is sampled from each distribution as introduced before.
+The results suggest that RIBBO has a strong generalizing ability to unseen function distributions.
+Due to the space limitation, the results and more details on cross-distribution generalization are deferred to Appendix D.
+The good generalization of RIBBO can be attributed to the paradigm of learning the entire algorithm, which can acquire general knowledge, such as exploration and exploitation trade-off from data, as observed in Chen et al.
+(2017).
 In contrast, such generalization may be limited if we learn surrogate models from data, because the function landscape inherent to surrogate models will contain only the knowledge of similar functions.
 (a) Visualization
-(b) Cross-dist performance
+(b) Generalization
 (c) Initial RTGs
@@
 Figure 3: (a) Visualization of the contour lines of $2$D Branin function and sampling points of RIBBO (red), Eagle Strategy (orange), and Random Search (gray), where the arrows represent the optimization trajectory of RIBBO.
-(b) Cross-distribution generalization by training the model on $4$ synthetic functions and testing on the remaining one, GriewankRosenbrock, which has different properties.
+(b) Generalization by training the model across all $24$ BBOB synthetic functions simultaneously.
+The results across functions with different output scaling are normalized to obtain the aggregate results.
 The legend shares with that of Figure 2.
 (c) Initial RTG $R_{0}$’s influence on performance.
-(d) RTG update strategy comparison between HRR and the naive strategy with various $R_{0}$.
+(d) RTG update strategy comparison between HRR and the naive strategy with various initial RTG $R_{0}$.
 4.4 Ablation Studies
@@
 Influence of Initial RTG Token $R_{0}$.
-By incorporating RTG tokens, RIBBO can attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$.
+By incorporating RTG tokens, RIBBO is able to attend to RTGs and generate optimization trajectories based on the specified initial RTG token $R_{0}$.
 To validate this, we examine the performance of RIBBO with different values of $R_{0}$, and the results are presented in Figure 3(c).
@@
 We can observe that the cumulative regrets of the generated query sequence do correlate with the specified RTG, indicating that RIBBO establishes the connection between regret and generation.
+We also conduct experiments to explore the effect of varying the immediate RTG $R_{t}$, and it is observed that setting to a value larger than 00 will decrease the performance and converge to a worse value.
+Please see Appendix E.
 Effectiveness of HRR.
 A key point of the inference procedure is how to update the value of RTG tokens at each iteration.
-To assess the effectiveness of the proposed strategy HRR in Eq.
+To assess the effectiveness of the proposed strategy HRR outlined in Eq.
 (5), we compare it with the naive strategy, that sets an initial RTG token $R_{0}$ and decreases it by the one-step regret after each iteration.
 The results are shown in Figure 3(d).
-The naive strategy displays distinct behaviors when $R_{0}=0$ or $R_{0}=5$.
-When $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially.
+The naive strategy displays distinct behaviors depending on the initial setting of $R_{0}$.
+Specifically, when $R_{0}=0$, i.e., the lower bound of regret, the model performs well initially.
 However, as the optimization progresses, the RTG tokens gradually decrease to negative values, leading to poor performance since negative RTGs are out-of-distribution values.
 Using $R_{0}=5$ compromises the initial performance, as the model may not select the most aggressive solutions with a high $R_{0}$.
-However, a higher $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage.
-The proposed HRR strategy performs the best across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens in Eq.
-(5) ensures the values remain meaningful and feasible.
+However, a higher initial $R_{0}$ yields better convergence value since it prevents out-of-distribution RTGs in later stage.
+The proposed HRR strategy consistently outperforms across the whole optimization stage, because setting the immediate RTG to 00 encourages the model to make the most advantageous decisions at every iteration, while hindsight relabeling of previous RTG tokens, as specified in Eq.
+(5), ensures that these values remain meaningful and feasible.
 Further Studies.
 We also study the effects of the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for $y$, the model size, and the sampled subsequence length $\tau$.
-Please see Appendix E.
-Generalization studies across the entire BBOB suite are shown in Appendix D, while the effects of transformations applied to the search space are detailed in Appendix F.
+Please see Appendix F.
+More visualizations illustrating the effects of random transformations on the search space are detailed in Appendix G.
 5 Conclusion
@@
 By incorporating RTG tokens into the optimization histories, RIBBO can automatically generate optimization trajectories satisfying the user-desired regret.
-Comprehensive experiments on BBOB, HPO and robot control problems show the versatility of RIBBO.
+Comprehensive experiments on BBOB synthetic functions, HPO and robot control problems show the versatility of RIBBO.
 This work is a preliminary attempt towards universal BBO, and we hope it can encourage more explorations in this direction.
 For example, we only consider the model training over continuous search space with the same dimensionality, and it would be interesting to explore heteroscedastic search space with different types of variables.
-Acknowledgments and Disclosure of Funding
-Use unnumbered first level headings for the acknowledgments.
-All acknowledgments go at the end of the paper before the list of references.
-Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work).
-More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
-Do not include this section in the anonymized submission, only in the final paper.
-You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
+A mathematical theoretical analysis of the in-context learning capabilities based on RTG tokens is also of interest.
 Appendix A Model Details
-We employ the commonly used GPT architecture [48] and the hyper-parameters are maintained consistently across the problems.
+We employ the commonly used GPT architecture Radford et al.
+(2018) and the hyper-parameters are maintained consistently across the problems.
 Details can be found in Table 1.
@@
 B.1 Benchmarks
-  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain.
+  • BBOB Elhara et al.
+(2019) is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain.
 This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties.
@@
   • SVM and XGBoost.
-HPO-B [2] is the most commonly used HPO benchmark and is grouped by search space id.
+HPO-B Arango et al.
+(2021) is the most commonly used HPO benchmark and is grouped by search space id.
 Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost.
@@
 We use the official implementations [^2] https://github.com/releaunifreiburg/HPO-B.
-  • Rover Trajectory Planning [16, 60] is a trajectory optimization task designed to emulate a rover navigation task.
+  • Rover Trajectory Planning Eriksson et al.
+(2019); Wang et al.
+(2018) is a trajectory optimization task designed to emulate a rover navigation task.
 The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, resulting in a total of $60$ parameters to optimize.
@@
 This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions.
-To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works [15, 59].
+To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works Elhara et al.
+(2019); Volpp et al.
+(2020).
 The training and test tasks are randomly sampled from the distribution.
@@
 The datasets are generated with several representatives of heuristic search, EA, and BO as behavior algorithms.
-We use the implementation in Open Source Vizier [52] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES.
+We use the implementation in Open Source Vizier Song et al.
+(2022) for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES.
 For Hill Climbing and Regularized Evolution, we provide a simple re-implementation.
-We use the implementation in BoTorch [8] for GP-EI.
+We use the implementation in BoTorch Balandat et al.
+(2020) for GP-EI.
 The details of these behavior algorithms are summarized as follows.
@@
 At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to generate $x_{next}$, which is then evaluated.
-If $f(x_{next})\textgreater f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
-  • Regularized Evolution [50] is an evolutionary algorithm with tournament selection and age-based replacement.
+If $f(x_{next})>f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
+  • Regularized Evolution Real et al.
+(2019) is an evolutionary algorithm with tournament selection and age-based replacement.
 We use a population size of $25$ and a tournament size of $5$.
@@
 The mutation operation uniformly selects one of the parameters and mutates it to a random value within the domain.
-  • Eagle Strategy [66] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies.
+  • Eagle Strategy Yang & Deb (2010) without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies.
 Each firefly emits light whose intensity corresponds to the objective value.
@@
 The position is updated based on the calculated weight.
-  • CMA-ES [26] is a popular evolutionary algorithm.
+  • CMA-ES Hansen (2016) is a popular evolutionary algorithm.
 At each iteration, candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
-  • GP-EI employs GP as the surrogate model and EI [30] as the acquisition function for BO.
+  • GP-EI employs GP as the surrogate model and EI Jones et al. (1998) as the acquisition function for BO.
+RIBBO is trained exclusively from offline datasets derived from various source tasks, which are sampled from the task distribution.
+Each dataset consists of several optimization histories, generated by running a behavior algorithm on a task.
+No additional assumptions are imposed to the data collection process or the behavior algorithms, thereby permitting the use of any behavior algorithm.
+Nevertheless, we can choose appropriate behavior algorithms to help the training.
+The model is trained on the datasets collected by the behavior algorithms, and the characteristics of these datasets can influence the model’s efficacy.
+It is advisable to employ well-performing and diverse behavior algorithms to create these datasets.
+If the datasets are from suboptimal behavior algorithms predominantly, there might be a decline in performance, even with the incorporation of RTG tokens.
+If the datasets are predominantly produced by behavior algorithms with homogeneous properties, the model may only be adept at addressing specific problems characterized by those properties.
+Conversely, if the behavior algorithms are diverse, the model can learn the strengths from various algorithms.
 B.3 Data
-For training, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data.
+For BBOB and rover problem, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data.
 Specifically, for the BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times, except for GP-EI, which is limited to $100$ function samples due to its high time cost.
-For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost.
-All behavior algorithms employ $500$ random seeds in the execution.
 For the rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times.
 For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
+For HPO-B, we use the “v3” meta-training/test splits provided by the authors, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost.
+All behavior algorithms employ $500$ random seeds to collect the training datasets.
 Appendix C Re-Implementation of OptFormer
-OptFormer [14] is a general optimization framework based on transformer [57] and provides an interface to learn policy or function prediction.
+OptFormer Chen et al.
+(2022) is a general optimization framework based on transformer Vaswani et al.
+(2017) and provides an interface to learn policy or function prediction.
 When provided with textual representations of historical data, OptFormer can determine the next query point $\bm{x}_{t}$, acting as a policy.
@@
 Appendix D Generalization
-To examine the generalization of our method to unseen function distributions, we conduct experiments to train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one.
+We conduct a series of experiments to examine the generalization of our method in this section.
+We train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one.
 The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
 Figure 5: Cross-distribution generalization by training on $4$ of $5$ chosen synthetic functions and testing on the remaining one.
-We also train the model across all training datasets from BBOB synthetic functions simultaneously, and evaluate on the same functions used in previous sections, specifically Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge.
-These functions serve as representatives of each category within the BBOB suite.
-The results are shown in Figure 6.
-RIBBO demonstrates relatively inferior performance compared to some behavior algorithms, which might be attributed to that the training dataset consists of excessive trajectories from diverse behavior algorithms across different functions, and complicates the process of algorithm identification solely through RTG tokens.
-Song et al.
-[53] had similar observation in Section 5.2 of their paper, indicating the significance of textual cues for $\bm{x}$ and the complete failure of the model when using anonymized inputs.
-However, RIBBO still surpasses both BC and BC Filter, emphasizing the effectiveness of RTG tokens.
-RIBBO employs RTG tokens to facilitate the identification of diverse algorithms across various functions and to generate sequences with specified regret.
-Further improvement could be achieved by integrating textual representations for $\bm{x}$ and metadata describing the problem, which may potentially improve the model’s ability to identify and generalize [54].
-This leaves as future work.
-Figure 6: Generalization by training across the entire BBOB suite.
-Appendix E Ablation Studies
+We also train the model across all training datasets from BBOB synthetic functions simultaneously, and normalize all the test functions to aggregate results across functions with different output scaling.
+The results are shown in Figure 6(a), and RIBBO has the best performance.
+(a) Aggregation on the entire BBOB
+(b) HRR with different $R_{t}$
+Appendix E Further Discussions about RTG Tokens
+We conduct more discussion about the setting of RTG tokens in this section.
+The immediate RTG tokens are designed to represent the performance based on cumulative regret over the future trajectories (i.e., $R_{i}=\sum_{t^{\prime}=i+1}^{t}(y^{*}-y_{t^{\prime}})$), rather than focusing solely on instantaneous regret $y^{*}-y_{t+1}$.
+Thus, setting the immediate RTG to 00 inherently accounts for future regret, thereby enabling the algorithm to autonomously trade-off the exploration and exploitation.
+We conduct experiments to examine the performance of RIBBO with different values of $R_{t}$.
+The results are shown in Figure 6(b).
+Setting to a value larger than 00 will decrease the performance and converge to a worse value, validating the effectiveness of setting the immediate RTG as 00.
+Sampling the next query point $\bm{x}_{t}$ is influenced by all the model inputs, including the immediate RTG token $\bm{x}_{t}$ and the previous history $(\bm{x}_{i},y_{i},R_{i})^{t-1}_{i=0}$.
+The immediate RTG token represents the goal that is intended to be achieved, while the historical data contains information about the problem.
+These elements are integrated to construct the inputs, influencing the resulting sampled point.
+Even if the RTG token is set to 00, implying a desire to sample the optimum, the short length of the history suggests limited knowledge about the problem, prompting the model to retain explorative behavior.
+As the history extends, suggesting a sufficient knowledge of the problem, the preference encoded by the RTG token shifts towards more greedy action.
+We conduct an experiment about the std of the output Gaussian head of $x$ of RIBBO using GriewankRosenbrock function and the results are shown in Table 2.
+It is clear that the std is very large in the early stage and becomes small later, which demonstrates the ability to trade-off the exploration and exploitation.
+Number of evaluations 1 30 60 90 120 150
+std of $\bm{x}$ 0.37 0.17 0.14 0.11 0.03 0.01
+Table 2: Std of the output Gaussian head of $x$.
+Appendix F Ablation Studies
 We provide further ablation studies to examine the influence of the key components and hyper-parameters in RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
@@
 The addition method integrates the values of each token into one, while the interleaving method addresses tokens sequentially.
-The results are shown in Figure 7(a).
+The results are shown in Figure 6(c).
 The concatenation method surpasses both the addition and interleaving methods, with the addition method showing the least efficacy.
@@
 No normalization uses the original function values directly.
-The results in Figure 7(b) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance.
+The results in Figure 6(d) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance.
 The choice of normalization method directly impacts the computation of RTG tokens and subsequently affects the generation of the desired regret.
@@
 Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating both training and inference.
-However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba and Grabocka [64] and Chen et al.
-[14].
+However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba & Grabocka (2021) and Chen et al.
+(2022).
 Therefore, we recommend using random normalization in practice.
@@
 The smaller model consists of $8$ layers, $4$ attention heads, and $128$-dimensional embedding space, while the larger model has $16$ layers, $12$ attention heads, and $384$-dimensional embedding space.
-The results are shown in Figure 7(c), indicating that the model size has a minimal impact on overall performance.
+The results are shown in Figure 6(e), indicating that the model size has a minimal impact on overall performance.
 Specifically, the smaller model, due to its limited capacity, shows a reduced performance, while the larger one, potentially more powerful, requires more training data, which can lead to a slight decrease in performance due to overfitting or inefficiency in learning from a limited dataset.
@@
 The process of subsequence sampling acts as a form of data augmentation, enhancing training efficiency.
-As shown in Figure 7(d), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance.
+As shown in Figure 6(f), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance.
 The computational complexity for causal transformer training and inference scales quadratically with the context length.
@@
 This highlights the trade-offs between computational complexity and performance.
-(a) Token aggregation
-(b) Normalization method
-(c) Model size
-(d) Subsequence length $\tau$
+(c) Token aggregation
+(d) Normalization method
+(e) Model size
+(f) Subsequence length $\tau$
 Figure 6: Ablation studies of token aggregation, normalization method, model size, and subsequence length $\tau$.
-Appendix F Visualization of Branin Function
+Appendix G Visualization of Branin Function
 In several of our experiments, such as BBOB, rover problems, and the visualization analysis of the Branin function, we implement a series of transformations to the search space.
@@
 Figure 7: Contour line visualization for samples drawn from the $2$D Branin function distribution using $4$ distinct random seeds.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The abstract and introduction clearly state the claims made, and the experimental results reflect the effectiveness of our method.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See the last paragraph of the paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: The paper does not include theoretical results.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: See Appendix A and B. The codes are provided in the supplementary.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The codes are provided in the supplementary and the scripts to reproduce experimental results are provided.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The experimental setting is presented in Section 4, and the full details are provided in Appendix A and B. The codes are provided in the supplementary.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We show error bars by the length of vertical bars in the figures.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: See Appendix A.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We have reviewed the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: There is no societal impact of the work performed.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: See Appendix B.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: The paper does not release new assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.